### v1

# LLM Meeting Decision Trees on Tabular Data

Abstract
Tabular data have been playing a vital role in diverse real-world fields, including healthcare, finance, etc. With the recent success of Large Language Models (LLMs), early explorations of extending LLMs to the domain of tabular data have been developed. Most of these LLM-based methods typically first serialize tabular data into natural language descriptions, and then tune LLMs or directly infer on these serialized data. However, these methods suffer from two key inherent issues: (i) data perspective: existing data serialization methods lack universal applicability for structured tabular data, and may pose privacy risks through direct textual exposure, and (ii) model perspective: LLM fine-tuning methods struggle with tabular data, and in-context learning scalability is bottle-necked by input length constraints (suitable for few-shot learning). This work explores a novel direction of integrating LLMs into tabular data throughough logical decision tree rules as intermediaries, proposes a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa. The proposed DeLTa avoids tabular data serialization, and can be applied to full data learning setting without LLM fine-tuning. Specifically, we leverage the reasoning ability of LLMs to redesign an improved rule given a set of decision tree rules. Furthermore, we provide a calibration method for original decision trees via new generated rule by LLM, which approximates the error correction vector to steer the original decision tree predictions in the direction of “errors” reducing. Finally, extensive experiments on diverse tabular benchmarks show that our method achieves state-of-the-art performance.
1 Introduction
Tabular data, typically organized in a structured table format within a relational database with rows and columns standing for the data samples and heterogeneous features (e.g., categorical and numerical features), is fundamental in various real-world fields, including healthcare hernandez2022synthetic , advertising frosch2010decade , finance assefa2020generating , etc. Given the heterogeneity of features borisov2022deep ; gorishniy2021revisiting , decision tree-based methods chen2016xgboost ; ke2017lightgbm ; prokhorenkova2018catboost were found to be particularly suitable for tabular data grinsztajn2022tree . While deep learning has led to breakthroughs in computer vision he2016deep and natural language processing devlin2018bert , decision tree-based methods still outperform the majority of existing deep tabular methods on tabular prediction tasks such as classification and regression. This superiority of decision tree-based methods is explained by the feature heterogeneity for tabular data, where neural networks struggle to learn the irregular target functions compared to decision tree-based methods grinsztajn2022tree .
Recently, LLMs have exhibited remarkable capabilities in natural language understanding and reasoning brown2020language ; achiam2023gpt , sparking growing interest in applying LLMs to structured tabular data tasks dinh2022lift ; hegselmann2023tabllm ; yan2024making ; wen2024supervised ; nam2024tabular ; han2024large ; yan2025small . Typically, most of the existing methods first serialize tabular samples into natural language descriptions, and then tune LLMs or directly infer via in-context learning on these serialized data. Despite the success of LLMs on text-based tasks, leveraging LLMs to empower tabular predictions remains a challenging task. Through an in-depth investigation of prior LLM-based approaches for tabular data, we identify two inherent characteristics hindering prediction:
(i) Data perspective: To bridge the modality gap between unstructured text and structured tabular data with heterogeneous features, most existing methods serialize tabular samples into text formats. For instance, TabLLM hegselmann2023tabllm converts a sample into listed feature descriptions such as “The [column name] is [feature value]”, and demonstrates that this template facilitates very-few-shot classification by leveraging the semantic priors of column names and values already encoded in LLMs. However, many real-world tabular datasets anonymize feature names using placeholder symbols for privacy wen2024supervised (e.g., finance), reducing the effectiveness of serialization methods that rely on semantic feature descriptors. In addition, LLMs tend to be less sensitive to numerical features fang2024large ; yan2024making . The predefined templates with inserted values often yield text formats that make LLMs struggle to understand the inherent interactions among different features. Beyond this, directly exposing the serialized tabular samples containing raw feature values to LLMs raises significant privacy concerns, especially in sensitive domains such as healthcare and finance, where data security is crucial yan2024protecting ; yao2024survey . These challenges make it difficult to design a universally suitable serialization template for tabular data.
Table 1: Most existing LLM-based tabular prediction methods require access to tabular samples and either train the LLMs or perform inference via in-context learning. Many of them are restricted to constrained settings such as few-shot or classification-only settings. The applicability of each scenario is determined on the experimental setups in the original papers. For instance, if a method primarily focuses on for few-shot learning, we mark the few-shot setting as applicable ( ); if it can also be extended to full-data training, we additionally mark the full-data setting as applicable ( ).
Methods No LLM access sample required No LLM training required Applied scenario
No pre-training No fine-tuning Full data Few shot Classification Regression
TabLLM (2023) hegselmann2023tabllm
LIFT (2022) dinh2022lift
TP-BERTa (2024) yan2024making
GTL (2024) wen2024supervised
SERSAL (2025) yan2025small ✔(Unsupervised) ✔(Binary classification)
P2T (2024) nam2024tabular
FeatLLM (2024) han2024large
DeLTa (Ours)
(ii) Model perspective: One straightforward way to adapt LLMs to tabular prediction tasks is fine-tuning LLMs’ parameters. For example, LIFT dinh2022lift investigated the fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional tabular prediction methods. Despite the effectiveness, fine-tuning LLMs remains a challenging task even with recent parameter-efficient fine-tuning methods han2024parameter , especially in structured tabular data dinh2022lift . Another research line employs in-context learning by adding few-shot example demonstrations to the prompts without training, which is evidenced by LLMs’ impressive capacity to learn from demonstrations included as part of the prompt brown2020language . But most of these methods are inherently limited by the input length constraints of LLMs, restricting their use to few-shot, or classification-only tasks. Therefore, LLMs still struggle to make satisfactory predictions for tabular data. We summarize these related works in Table 1.
In this paper, we explore a novel direction of integrating LLMs into tabular data via logical decision tree rules loh2011classification as intermediaries. We propose DeLTa, a decision tree enhancer with LLM-derived rule for tabular prediction. Specifically, we first leverage the reasoning ability of LLMs to redesign an improved rule given a decision tree rule set, enforcing greater coherence among samples falling into the same leaf node. The newly generated rule is used to approximate a sample-specific error correction vector to calibrate the prediction of original decision trees in the direction of “errors” reducing. The proposed DeLTa could well solve the aforementioned challenges: Solving the data issue: Unlike serialization methods that convert each sample into unnatural text formats, decision tree rules are composed of simple comparisons between feature values and thresholds, forming logical, interpretable structures that can be naturally expressed in text without relying on semantic columns names. In addition, decision tree rules represent global feature space partitioning rule rather than individual samples, which helps mitigate privacy concerns by avoiding exposure of sample-level information. Solving the model issue: Moreover, the powerful reasoning ability of LLMs can be leveraged to redesign decision tree rules and help trees with aggregating their decisions, rather than directly using LLMs to generate label predictions. Notably, DeLTa avoids serialize tabular data into natural language format, and does not require additional domain-specific expertise or semantic information, such as explicit feature names and detailed task background knowledge. Furthermore, DeLTa can be applied in full data learning setting without LLM fine-tuning.
The contributions of this paper include: 1) We investigated the prior LLM-based methods for tabular prediction and explore a novel direction of integrating LLMs into tabular data via refining the decision tree rules, without directly accessing to the data itself, addressing the inherent issue w.r.t. data perspective and model perspective. 2) We propose a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa, which utilizes LLMs to refine a set of decision tree rules derived on decision trees trained on multiple train subsets, where the newly generated rule is used to infer the sample-wise error correction vector to calibrate the output of original decision trees. 3) We conduct extensive experiments on various tabular benchmarks and competing benchmark algorithms, and comprehensive results along with analysis and visualizations demonstrate our effectiveness.
2 Related work
Machine learning for tabular prediction. The development of effective algorithms for predictive modeling on tabular data has been a longstanding research topic. In the early days, decision tree-based methods (e.g. XGBoost chen2016xgboost , CatBoost prokhorenkova2018catboost ) were found to be particularly suitable for tabular data. More recently, inspired by the success of of deep learning in computer vision (CV) he2016deep and natural language processing (NLP) devlin2018bert , numerous methods have been proposed for tabular data to accomplish tabular prediction tasks. These works mainly include MLP-like models wang2021dcn , attention-based architectures song2019autoint ; gorishniy2021revisiting ; somepalli2021saint , and retrieval-augmented architectures ye2025revisiting . Among these works, DCN V2 wang2021dcn is an architecture that consists of an MLP-like module and a feature crossing module; AutoInt song2019autoint leveraged the Transformer architecture to capture inter-column correlations; FT-Transformer gorishniy2021revisiting further enhanced AutoInt’s performance through improved token embeddings; ModernNCA ye2025revisiting makes predictions based on the relationships with neighbors in a learned embedding space. Recently, another line of research has tried to use additional information outside target dataset to enhance tabular data prediction. XTab zhu2023xtab pretrains Transformer on a variety of datasets for cross-table pretraining. TabPFN hollmann2025accurate ; hollmann2023tabpfn , which is pretrained on a large set of synthetic datasets, serves as a foundation model for small to medium-sized tabular data.
Large language models for tabular prediction. Motivated by the impressive success of LLMs brown2020language ; achiam2023gpt , another promising study has attempted to harness the rich prior knowledge encapsulated by LLMs in tabular prediction tasks. LIFT dinh2022lift investigated the performance of fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional solutions. Extending this line of research, TabLLM hegselmann2023tabllm employed T0 sanh2022multitask as the base LLM and demonstrated competitive performance of fine-tuned LLMs in very few-shot scenarios. Additionally, recent efforts have focused on pre-training LLMs over diverse tabular datasets from different domains. Among these, TP-BERTa yan2024making introduces a tabular-specific tokenization scheme, enabling a single pre-trained language model to generalize across multiple tabular datasets after further fine-tuning. GTL wen2024supervised further promotes comprehensive instruction-following capabilities for both zero-shot and in-context learning with a limited number of examples. Despite their promising results, these approaches typically require training or fine-tuning LLMs. Instead of directly fine-tuning LLMs, P2T nam2024tabular leverages unlabeled data correlated with target data expressed in natural language form and prompts LLMs for few-shot semi-supervised learning. Recently, Summary manikandan2023language proposed a boosting framework that treats LLMs as weak learners for tabular prediction, particularly for tasks involving small numbers of data points. FeatLLM han2024large prompts LLMs with serialized training examples to generate new features for few-shot classification tasks. SERSAL yan2025small explores to use synergy learning with FT-Transformer to enhance the noisy annotations generated by LLMs for binary classification tasks in unsupervised manner. This method also needs to fine-tune the LLMs. Going beyond tabular prediction, Nam et al. nam2024optimized leverages LLMs to optimize the feature generator with tree rules in the field of feature engineering zhang2023openfe , which we do not consider as a close related work to ours as the primary goal, motivation, and methodology are different.
Although prior works demonstrate that LLMs can make predictions for tabular tasks, the majority of these approaches rely on first converting tabular data into natural language descriptions, i.e., serialization, which often produces unnatural texts that differ from how humans might describe the data manikandan2023language . And it is also challenging to design such suitable template. In addition, these methods suffer from two limitations that either need to train LLMs or are restricted to constrained settings such as few-shot or classification-only settings. Table 1 summarizes these related works. The above limitations raise a fundamental question: can we harness the capabilities of LLMs without fine-tuning and without serializing tabular data, to improve full-data tabular prediction tasks across both classification and regression?
3 Preliminaries
Problem formulation. Denote a tabular dataset $\mathcal{D}\!=\!\{(x_{i},y_{i})\}_{i=1}^{N}$ as a collection of $N$ samples, where $(x_{i},y_{i})$ is the $i$-th data pair with $x_{i}\in\mathbb{X}$ representing input features and $y_{i}\in\mathbb{Y}$ the corresponding label. Concretely, we have $x_{i}\!=\!(x_{i}^{(num)},x_{i}^{(cat)})$ where $x_{i}^{(num)}$ and $x_{i}^{(cat)}$ represent the numerical and categorical features respectively. We consider supervised tabular prediction tasks: binary classification $\mathbb{Y}\!=\!\{0,1\}$, multiclass classification $\mathbb{Y}\!=\!\{1,...,c\}$ and regression $\mathbb{Y}\!=\!\mathbb{R}$. For data splits, $\mathcal{D}_{train}$ denotes training set for model training, $\mathcal{D}_{val}$ validation set for early stopping and hyperparameter tuning, and $\mathcal{D}_{test}$ test set for final evaluation. The goal is to obtain an accurate model $G:\mathbb{X}\to\mathbb{Y}$ trained on $\mathcal{D}_{train}$, that minimizes the expected loss $\mathbb{E}[\mathcal{L}(G(x),y)]$. Here, $\mathcal{L}$ is a smooth loss function (e.g., mean squared error or cross-entropy) and $G$ can be any tabular predictive model.
Decision trees (DTs). A decision tree (DT) loh2011classification is a hierarchical model that recursively partitions the input feature space into disjoint regions, i.e., leaf nodes, through a series of axis-aligned splits. For a given partitioning rule $r$, the prediction function $f$ can be formally expressed as:
$f(x|\mathcal{D}_{train},r)=\sum_{l=1}^{\text{Node}(r)}\lambda_{l}\cdot\mathbb{I}(x\in L_{l}(r)),$ (1)
where $\text{Node}(r)$ denotes the number of leaf nodes in the tree, $L_{l}(r)$ denotes the $l$-th leaf node containing a set of $x$ from $\mathcal{D}_{train}$, $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r)$ based on rule $r$, and $\lambda_{l}:\mathbb{X}\to\mathbb{Y}$ represents leaf-specific prediction values. Typically, $\lambda_{l}$ will be the average of corresponding labels for the leaf node in regression tasks, or the empirical class distribution in classification tasks. The structure of the tree is governed by a feature index vector $\ell$ and a threshold vector $\tau$, which together define the decision path containing a series of internal nodes from the root to each leaf node. Specifically, decision tree is constructed by recursively applying binary tests of the form $x^{\ell_{j}}\leq\tau_{j}$ at internal nodes, where $j$ indexes internal node along the root-to-leaf path, $\ell_{j}$ specifies feature index for $j$-th internal node split, $x^{\ell_{j}}$ denotes the corresponding feature value, and $\tau_{j}$ defines corresponding splitting threshold. In this work, we adopt CART loh2011classification as our decision tree implementation due to two primary reasons: (i) it offers high interpretability and can be expressed using a simple if-else syntax, (ii) tree-based models, often ensembles of simple decision trees like CART, outperform deep learning approaches in numerous tabular prediction tasks grinsztajn2022tree ; gorishniy2021revisiting .
4 Proposed method: DeLTa
We propose a novel approach, Decision Tree Enhancer with LLM-derived Rule for Tabular Prediction (DeLTa), understanding the underlying rule logic encoded by decision trees and generating improved rules for tabular prediction via LLM, without requiring any fine-tuning of the LLM. This strategy introduces a fundamentally distinct interface for extending LLM capabilities to tabular tasks, departing from the majority of conventional data-to-text serialization pipelines that enforce LLMs to understand the serialized tabular data. In the following, we will elaborate on LLM-based decision tree rules refinement in Section 4.1; then, we will provide the calibration method for original decision trees via new generated rule by LLM in Section 4.2; we also give the overall implementation of ours in Section 4.3. An overview of the proposed framework is depicted in Fig. 1.
4.1 LLM-based decision tree rules refinement
The decision tree rules refinement procedure consists of the following stages: (i) decision tree rules initialization, that constructs a diverse decision tree rule set, and (ii) rule understanding via LLM, that leverages the LLM to understand and refine the original rule set to obtain a new rule.
Decision tree rules initialization. To mitigate the risk of overfitting for a single decision tree expert ho1998random , the classical Random Forest breiman2001random algorithm trains diverse decision tree experts $\{f_{k}\}_{k=1}^{K}$ on different subsets of $\mathcal{D}_{train}$, and uses the ensemble of $\{f_{k}\}_{k=1}^{K}$ as prediction:
$F(x)=\frac{1}{K}\sum_{k=1}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k}),$ (2)
where $F(x)$ denotes the label prediction produced by the Random Forest, $K$ is the number of experts, $\mathcal{D}_{train}^{k}$ denotes the subset, $r_{k}$ denotes the rule for expert $f_{k}$ derived from Eq. 1, and =$\mathcal{R}=\{r_{k}\}_{k=1}^{K}$ denotes a source decision tree rule set. In terms of decision tree rule, each rule represents a logical comparison between features and thresholds displayed in semantically rich syntax, which thus can be used to partition the feature space into disjoint regions (i.e., leaf nodes) with high interpretability. A well understanding of these rules could generate a better feature space partitioning method, thereby promoting statistical coherence among samples assigned to the same leaf node. Although Random Forest ensembles the outputs from all rules, the inherent relationships and interactions among rules in $\mathcal{R}$ are ignored. With the development of $K$, analyzing these independent rules is gradually becoming more and more difficult, let alone utilizing these rules to partition feature space. To this end, we propose to leverage LLMs to analyze and summarize the rule set $\mathcal{R}$ into a refined rule due to the powerful logical reasoning ability of LLM.
Figure 1: The DeLTa framework. As shown in the main objective, we calibrate the output of original decision tree experts $F(x)$ in the direction of “errors” reducing. Subfig (a) describes the process of refining decision tree rules with LLM, and subfig (b) details the refined rule-guided error correction for decision trees.
Rule understanding via LLM. Our objective is to use an LLM to generate a new rule $r^{*}$ given original rule set $\mathcal{R}$. Specifically, we construct a prompt $p$ for the LLM that guides it to understand the diverse decision tree rule set $\mathcal{R}$ and synthesize the redesigned rule from the aggregated diverse knowledge. To effectively leverage the reasoning ability of LLM, the prompt $p$ (please see Appendix A.4) is designed to include: meta information $p_{meta}$ that describes the task objective, the extracted decision tree rule description $p_{rule}$ that contains $\mathcal{R}$, and the requirement $p_{requirement}$ for rule refinement. The formulation of generating $r^{*}$ is given by:
$r^{*}=LLM(p)=LLM(p_{meta}\oplus p_{rule}\oplus p_{requirement}),$ (3)
where $\oplus$ is the concatenation of each prompt, and this process is achieved through querying the LLM, without requiring any fine-tuning. Notably, our approach does not require additional domain-specific expertise or semantic information, such as explicit feature names or detailed task background knowledge. This makes it applicable to scenarios where such information is unavailable or incomplete. Instead of relying on the intrinsic domain knowledge of tabular tasks, such as those in healthcare or finance, we frame the problem purely as an domain knowledge-agnostic machine learning task. Our method leverages LLMs to reason decision tree rule set $\mathcal{R}$ into an rule $r^{*}$, rather than directly inferring from serialized tabular data. To verify that LLM could generate a better rule, where samples grouped within the same leaf node exhibit greater statistical similarity, we compute the intra-node sample distance over all leaf nodes partitioned by $r^{*}$, and observe that the distance of $r^{*}$ is lower than original rule $r\in\mathcal{R}$, as illustrated in Fig. 2. Next, we will elaborate on how to leverage the new rule $r^{*}$ generated by LLM to enhance tabular predictions. [Uncaptioned image] Figure 2: Average intra-node distance comparison.
4.2 Refined rule-guided decision tree error correction
The refined rule $r^{*}$ is summarized over the rule set $\mathcal{R}$, and thus provides a potentially better strategy that partitions the feature space of $\mathcal{D}_{train}$ into disjoint leaf nodes $\{L_{l}(r^{*})\}_{l=1}^{\text{Node}(r^{*})}$, with each leaf node containing a subset of $x$ from $\mathcal{D}_{train}$ and $\text{Node}(r^{*})$ denoting the number of leaf nodes. To use $r^{*}$ for making predictions, the simplest strategies include: (1) treating it as a standalone decision tree by assigning the input $x$ to a specific leaf node, or (2) appending it to the existing ensemble of decision trees $F(x)$. However, the former may be too isolated, as it does not leverage the predictive power of the existing ensemble, while the latter may be too diluted, as it does not fully utilize the refined guidance provided by LLMs. This work introduces a novel way to leverage the power of $r^{*}$ together with $F(x)$ inspired by the intuition that estimating the difference between predictions and labels, i.e., the residual errors is often easier than directly predicting the ground-truth labels zhang2005boosting ; he2016deep . Intuitively, we first use $r^{*}$ to partition data samples into its leaf nodes, expecting the samples in the same node are similar to each other. Motivated by the famous Gradient Boosting algorithm friedman2001greedy ; friedman2002stochastic , for the samples in each node, we learn a specific mapping function that predicts residual errors of $F(x)$ for the samples in that node. These learned functions are subsequently used to adjust the predictions made by $F(x)$, leading to improved accuracy. Since $r^{*}$ is designed to group similar samples effectively, the residuals within each leaf node are expected to be more structured and thus easier to predict.
Formally, we introduce the definition of the direction of prediction “errors” increasing and reducing, and the definition of a new set derived from $\mathcal{D}_{train}$ used for fitting the mapping functions as follows.
Definition 1. Prediction “errors” increasing and reducing. For a given training sample $x$, we approximate the "error" using the gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$, where $\mathcal{L}$ is the loss function discussed in Section 3. Intuitively, shifting $F(x)$ in the direction of this gradient (“Errors” Increasing) would likely worsen the model’s performance, analogous to how traditional gradients indicate a direction in parameter space that increases the loss ruder2016overview . Therefore, it is natural that we should calibrate $F(x)$ in the direction of negative gradient $-\nabla_{F(x)}\mathcal{L}(F(x),y)$ (“Errors” Reducing).
Definition 2. Gradient set $\mathcal{D}_{train}^{\nabla}$. $\mathcal{D}^{\nabla}_{train}=\{(x,-\nabla_{F(x)}\mathcal{L}(F(x),y))|(x,y)\in\mathcal{D}_{train}\}$ is derived from train set $\mathcal{D}_{train}$ and stores training sample features and their corresponding negative gradients of the loss function with respect to the output of $F(x)$. The only difference between $\mathcal{D}_{train}$ and $\mathcal{D}_{train}^{\nabla}$ lies in the labels, which are replaced by the negative gradients in $\mathcal{D}_{train}^{\nabla}$, while $r^{*}$ induces the same leaf node partitioning of the feature space for both datasets.
After obtaining the Gradient Set $\mathcal{D}_{train}^{\nabla}$, we apply a learnable Gradient Net $\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ to model the mapping between $x$ and the corresponding negative gradient within $\mathcal{D}_{train}^{\nabla}$, where each leaf node $L_{l}(r^{*})$ corresponds to one specific mapping function $\phi_{l}(\cdot;\theta_{l}):x\to-\nabla_{F(x)}\mathcal{L}(F(x),y)$. The training process for $\phi_{l}$ is as follow:
$\min_{\theta_{l}}\sum_{x\in L_{l}(r^{*})}\|\phi_{l}(x;\theta_{l})-(-\nabla_{F(x)}\mathcal{L}(F(x),y))\|_{2}^{2},$ (4)
where $\phi_{l}$ could be easily implemented by conventional machine learning model, such as CART loh2011classification , and $\phi_{l}$ over different leaf nodes are trained separately.
At inference time, given $x$ from $\mathcal{D}_{test}$, the prediction of Gradient Net $\phi$ could be produced as a sample-specific error correction vector $\Delta_{x}$. The final prediction $F^{*}(x)$ could be obtained by adding $\Delta_{x}$ to $F(x)$, that moves $F(x)$ in the direction of negative gradient (“Errors” Reducing) by one step:
$\Delta_{x}=\eta*\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})=\eta*\sum_{l=1}^{\text{Node}(r^{*})}\phi_{l}(x;\theta_{l})\cdot\mathbb{I}(x\in L_{l}(r^{*})),$ (5)
$F^{*}(x)$ $=F(x)+\Delta_{x},$ (6)
where $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r^{*})$ based on rule $r^{*}$, $\eta\in\mathbb{R}_{+}$ is the hyperparameter similar to learning rate controlling the step size, and $\Delta_{x}$ approximates $-\eta*\nabla_{F(x)}\mathcal{L}(F(x),y)$. We now explain the proposed error correction strategy by giving a closer look at Eq. 6.
Proposition 1
Let $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ denote the expected loss, where $F(x)=\frac{1}{K}\sum_{k}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k})$ and each $f_{k}$ corresponds to a decision tree rule $r_{k}$ from the expert-derived rule set =$\mathcal{R}=\{r_{k}\}_{k=1}^{K}$. Given a prompt $p$ that contains $\mathcal{R}$, the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ could be decreased by querying the LLM with $p$ to generate a refined rule $r^{*}$, which enables approximation of a sample-specific error correction vector $\Delta_{x}$ to guide prediction $F(x)$ in the direction of “errors” reducing.
The proof of Proposition 1 is provided in Appendix A.5. To reduce the expected loss, we can approximate and apply negative gradient-based error correction vector $\Delta_{x}$ to $F(x)$.
4.3 Framework overview and discussion
Framework overview. In terms of training stage, the procedure begins by training a set of different decision tree experts (using CART for each expert) and obtaining the ensemble predictions based on Eq. 2. We then extract the corresponding decision tree rule set $\mathcal{R}$. The set of rules is used to construct a prompt $p$ for LLM, which returns an improved rule $r^{*}$ (Eq. 3). Subsequently, we compute the negative gradient for each training sample to construct the Gradient Set $\mathcal{D}_{train}^{\nabla}$, and then train the Gradient Net $\phi$ based on $\mathcal{D}_{train}^{\nabla}$ and $r^{*}$ (Eq. 4). During the test stage, we can achieve the preliminary prediction result $F(x)$ for each test sample $x$ based on the random forest; and also compute the sample-specific error correction vector $\Delta_{x}$ by feeding the test sample to Gradient Net. Now the final prediction can be corrected using Eq. 6. The complete training and inference pipeline is summarized in Algorithm 1 of Appendix A.3.
Discussion. To sum up, fed with the expert-derived rules $\mathcal{R}$, LLM is prompted to summarize multiple rules and redesign a new rule $r^{*}$, which thus considers the inherent relationships and interactions among multiple independent rules. Benefiting from the rule-based prompt, we neither need to fine-tune LLM nor expose the serialized samples into LLM. Due to the powerful reasoning ability of LLM, the refined rule $r^{*}$ that summarizes multiple rules is more effective and also convenient for us to use. For example, we can not only use the rule to make decision but also to partition the feature space of training set. The latter means that we can build a set of disjoint leaf nodes, where samples grouped within the same leaf node exhibit greater statistical similarity and thus can be used to train network. Here, we use the samples in the same leaf node to model the mapping between sample and the corresponding negative gradient, which is usually easier than directly fitting the relationship between samples and the labels. Our proposed DeLTa enables a principled integration of expert knowledge and LLM-based rule synthesis for enhanced performance, which paves a new way for LLM-based tabular data learning.
5 Experiments
Datasets. We consider a variety of supervised tabular prediction tasks with heterogeneous features, including binary classification, multiclass classification, and regression. Specifically, the tabular datasets include: Blood (BL) blood_transfusion_service_center_176 , Credit (CR) statlog_(german_credit_data)_144 , Car bohanec1988knowledge , Bank (BA) Moro2014ADA , Adult (AD) kohavi1996scaling , Jannis (JA) grinsztajn2022tree , Cpu_act (CP) grinsztajn2022tree , Credit_reg (CRR) grinsztajn2022tree , California_housing (CA) pace1997sparse , House_16H (HO) grinsztajn2022tree , Fried (FR) breiman1996bagging , Diomand (DI) chen2023trompt . The dataset properties and data pre-processing details are summarized in Appendix A.1. Following previous studies gorishniy2021revisiting , we use Accuracy (higher is better) to evaluate binary and multiclass classification tasks, Normalized Root Mean Squared Error (NRMSE) (lower is better) to evaluate the regression tasks.
Baselines and implementation details. We conduct a comparative analysis between DeLTa and other prominent methods in the field of tabular prediction. Specifically, we include LLM-based methods such as TabLLM hegselmann2023tabllm , LIFT dinh2022lift , TP-BERTa yan2024making , GTL wen2024supervised , P2T nam2024tabular , FeatLLM han2024large . For LIFT, we employ two versions, i.e., fine-tuning LLMs (LIFT) and in-context learning (LIFT-ICL) versions, as provided in their original paper. In addition, we compare DeLTa with non LLM-based methods, including KNN cover1967nearest , CART loh2011classification , MLP lecun2015deep , Random Forest breiman2001random , XGBoost chen2016xgboost , CatBoost prokhorenkova2018catboost , FT-Transformer gorishniy2021revisiting , TabPFN hollmann2025accurate , ModernNCA ye2025revisiting . More details about the baseline models can be found in Appendix A.2. Implementation details of DeLTa including hyper-parameters are provided in Appendix A.3. For LLM usage, DeLTa adopts GPT-4o as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix A.6 for more results). To prevent LLMs from generating task-irrelevant output, we query LLM 10 times by default and use the average. In replicating baselines, GPT-4o is utilized for in-context learning. For baselines require LLM fine-tuning, we employ the LLMs used in their original paper: TabLLM uses T0 sanh2022multitask ; we use the API provided by OpenAI to perform black-box GPT-3.5 fine-tuning for LIFT, where the fine-tuning method is suggested by the original paper; TP-BERTa employs RoBERTa liu2019roberta ; GTL uses the 13B version of LLaMA 2 touvron2023llama . The experiments are run on NVIDIA A100-PCIE-40GB GPU. To reduce the effect of randomness, the reported performance is averaged over 5 independent runs.
5.1 Main Results
Comparison with LLM-based methods. We evaluated the performance of DeLTa against mainstream LLM-based baseline methods on tabular classification tasks, as illustrated in Table 2. In addition to standard prediction tasks under full-data regimes, we also provide results under low-data regimes (few-shot learning), as some of these LLM-based baseline methods are constrained by the limited input length of LLMs and are thus restricted to low-data settings. Following prior work han2024large , we simulate the few-shot setting by randomly sampling a fixed number of training examples, where the number of shots denotes the total number of selected samples. Considering some of the LLM-based baseline methods are applicable to regression tasks, we further evaluate DeLTa on regression under both full-data and low-data regimes settings, as illustrated in Fig. 3. Overall, DeLTa achieves the highest average performance in all settings, including classification and regression, in both full and low-data regimes. While DeLTa is primarily designed for full-data tabular prediction, our results demonstrate that it also performs competitively in low-data scenarios. Furthermore, compared to existing LLM-based tabular prediction methods, the proposed DeLTa requires no additional requirements for LLMs as illustrated in Table 2. It also aligns with our claims that we have solved the two inherent issues for prior LLM-based approaches: (i) the proposed DeLTa avoids reliance on column names and tabular data serialization, and thus will protect sample-level privacy information, and (ii) rather than directly using LLMs to generate label predictions, DeLTa utilizes the powerful reasoning ability of LLMs to enhance decision trees, enabling effective tabular prediction in both full-data and low-data regimes without requiring LLM fine-tuning.
Table 2: Test Accuracy ($\uparrow$) performance of DeLTa and LLM-based baseline methods on classification tasks. “# Num” indicates the number of training samples. “Tab” indicates whether the LLM requires access to tabular samples. “Extra” indicates whether the method needs to use extra data samples outside the training set. “Feat” indicates whether the feature names are needed. “Tune” indicates whether LLM fine-tuning is needed.
# Num Method (LLM used) Additional requirements Datasets
Tab Extra Feat Tune BL $\uparrow$ CR $\uparrow$ Car $\uparrow$ BA $\uparrow$ AD $\uparrow$ JA $\uparrow$ Average $\uparrow$
All TabLLM (T0) 0.761 0.737 0.845 0.906 0.864 0.691 0.801
All LIFT (GPT-3.5) 0.689 0.691 0.729 0.825 0.81 0.647 0.732
All TP-BERTa (RoBERTa) 0.761 0.730 0.826 0.916 0.844 0.659 0.789
All GTL (LLaMA2) N/A N/A N/A N/A N/A N/A N/A
All LIFT-ICL (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
All P2T (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
All FeatLLM (GPT-4o) 0.768 0.701 0.589 0.887 0.842 0.540 0.721
All DeLTa (GPT-4o) 0.829 0.783 0.836 0.908 0.868 0.705 0.822
256 TabLLM (T0) 0.344 0.690 0.471 0.785 0.787 0.373 0.575
256 LIFT (GPT-3.5) 0.651 0.682 0.431 0.724 0.773 0.398 0.610
256 TP-BERTa (RoBERTa) 0.721 0.677 0.408 0.853 0.772 0.464 0.649
256 GTL (LLaMA2) N/A N/A N/A N/A N/A N/A N/A
256 LIFT-ICL (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
256 P2T (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
256 FeatLLM (GPT-4o) 0.729 0.658 0.459 0.740 0.772 0.473 0.639
256 DeLTa (GPT-4o) 0.732 0.713 0.736 0.767 0.778 0.476 0.700
64 TabLLM (T0) 0.245 0.673 0.386 0.602 0.785 0.268 0.493
64 LIFT (GPT-3.5) 0.643 0.643 0.500 0.728 0.748 0.430 0.615
64 TP-BERTa (RoBERTa) 0.648 0.646 0.419 0.840 0.761 0.414 0.621
64 GTL (LLaMA2) 0.613 0.638 0.512 0.719 0.750 0.414 0.608
64 LIFT-ICL (GPT-4o) 0.348 0.538 0.326 0.469 0.613 0.251 0.424
64 P2T (GPT-4o) 0.368 0.542 0.334 0.569 0.634 0.243 0.448
64 FeatLLM (GPT-4o) 0.697 0.643 0.508 0.723 0.749 0.450 0.628
64 DeLTa (GPT-4o) 0.732 0.663 0.615 0.733 0.758 0.457 0.660
Figure 3: Test NRMSE ($\downarrow$) performance of DeLTa and LLM-based baseline methods on regression tasks.
Table 3: Runtime in seconds of DeLTa and other LLM-based baseline methods for the training and inference phase, conducted on Adult dataset. “# Num” indicates number of training samples. †These methods require querying LLMs for each test sample. * These methods do not require querying LLMs at inference phase.
# Num Stage TabLLM† LIFT† TP-BERTa† GTL† LIFT-ICL† P2T† FeatLLM* DeLTa* (Ours)
All Train 177371.37 153689.14 1319.35 N/A N/A N/A 1231.22 35.20
Inference 179.21 90149.48 1.58 N/A N/A N/A 0.14 0.09
64 Train 288.61 191.75 316.40 N/A N/A N/A 1047.55 23.09
Inference 170.42 89684.76 5.01 153501.02 176726.13 200258.24 0.14 0.08
Computational efficiency. We compare the computational cost between ours and LLM-based methods in Table 3. The results show that our proposed DeLTa is computationally efficient in both training and inference stages. This efficiency stems from key properties: (i) DeLTa utilizes the reason ability of LLM without fine-tuning LLM, (ii) DeLTa avoids querying LLMs to generate predictions for individual samples. Instead, it only needs to query LLM via API to generate one refined decision tree rule for one dataset, enabling significantly more efficient use of LLM resources.
Comparison with conventional methods. We further compare the performance of DeLTa and other conventional tabular models on full data regimes averaged over all datasets, as illustrated in Fig. 4. Full results are provided in Appendix A.8. The results show that DeLTa also performs well compared with conventional tabular models on both classification and regression tasks.
Figure 4: Test performance of DeLTa and non-LLM baseline methods on classification and regression tasks. Here, “RF” denotes Random Forest, “FT-T” denotes FT-Transformer, “MNCA” denotes ModernNCA.
5.2 Further analysis
Ablation study. We further conduct ablation study to demonstrate the effectiveness of key components of DeLTa, as illustrated in Table 4, by comparing DeLTa against three variants: (i) DeLTa w/o “RR”, that replaces the LLM generated $r^{*}$ used in Gradient Net $\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ with $r\in\mathcal{R}$ derived from Random Forest $F(x)$, (ii) DeLTa w/o “EC”, that directly obtains the label prediction for test $x$ by utilizing $f(x|\mathcal{D}_{train},r^{*})$ (Eq. 1) with the refined rule $r^{*}$, (iii) Random Forest, which makes decision with $F(x)$ in Eq. 2. The results show that the decision tree rule refinement via LLM is important, and the error correction strategy could further enhance tabular prediction performance.
Table 4: Ablation study on the different components in DeLTa; see full results in Appendix A.10. “RR” denotes the rule refinement process, and “EC” denotes the error correction process.
Variants CR $\uparrow$ CA $\downarrow$
DeLTa (Ours) 0.7829 0.3507
w/o “RR” 0.7257 0.5443
w/o “EC” 0.7657 0.4016
Random Forest 0.7371 0.5057
Additional results. Fig. 5 shows the label predictions of ours: $F(x)+\Delta_{x}$ and ours w/o $\Delta_{x}$: $F(x)$. We can observe that two categories of samples are overlapped in some regions. $F(x)$ struggles to distinguish them, but DeLTa could correct the prediction errors of $F(x)$ to enhance the prediction for such complex data patterns.
Figure 5: Visualization of label prediction of DeLTa w/ and w/o error correction vector $\Delta_{x}$ on BA dataset.
We incorporate the sensitivity analysis for hyperparameters in Appendix A.11. Additional insights and analysis regarding the LLM-refined rule can be found in Appendix A.12.
6 Conclusion
In this paper, we propose DeLTa, a decision tree enhancer with LLM-derived rule for tabular prediction, to solve the two challenges w.r.t. data perspective and model perspective in prior LLM-based approaches for tabular prediction. DeLTa avoids reliance on feature names and tabular data serialization, and it utilizes the powerful reasoning ability of LLMs to enhance decision trees, enabling effective tabular prediction in both full-data and low-data regimes without requiring LLM fine-tuning. The empirical results demonstrated the effectiveness of DeLTa for tabular prediction tasks. Our work can shed some light on developing better algorithms for similar tasks.