### v2

# LLM Meeting Decision Trees on Tabular Data

Abstract
Tabular data have been playing a vital role in diverse real-world fields, including healthcare, finance, etc. With the recent success of Large Language Models (LLMs), early explorations of extending LLMs to the domain of tabular data have been developed. Most of these LLM-based methods typically first serialize tabular data into natural language descriptions, and then tune LLMs or directly infer on these serialized data. However, these methods suffer from two key inherent issues: (i) data perspective: existing data serialization methods lack universal applicability for structured tabular data, and may pose privacy risks through direct textual exposure, and (ii) model perspective: LLM fine-tuning methods struggle with tabular data, and in-context learning scalability is bottle-necked by input length constraints (suitable for few-shot learning). This work explores a novel direction of integrating LLMs into tabular data through logical decision tree rules as intermediaries, proposing a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa. The proposed DeLTa avoids tabular data serialization, and can be applied to full data learning setting without LLM fine-tuning. Specifically, we leverage the reasoning ability of LLMs to redesign an improved rule given a set of decision tree rules. Furthermore, we provide a calibration method for original decision trees via new generated rule by LLM, which approximates the error correction vector to steer the original decision tree predictions in the direction of “errors” reducing. Finally, extensive experiments on diverse tabular benchmarks show that our method achieves state-of-the-art performance. The source code is available at https://github.com/HangtingYe/DeLTa.
1 Introduction
Tabular data, typically organized in a structured table format within a relational database with rows and columns standing for the data samples and heterogeneous features (e.g., categorical and numerical features), is fundamental in various real-world fields, including healthcare hernandez2022synthetic, advertising frosch2010decade, finance assefa2020generating, etc. Given the heterogeneity of features borisov2022deep; gorishniy2021revisiting; ye2024ptarl; ye2025drl, decision tree-based methods chen2016xgboost; ke2017lightgbm; prokhorenkova2018catboost were found to be particularly suitable for tabular data grinsztajn2022tree; fan2020autofs; fan2021interactive. While deep learning has led to breakthroughs in computer vision he2016deep and natural language processing devlin2018bert, decision tree-based methods still outperform the majority of existing deep tabular methods on tabular prediction tasks such as classification and regression. This superiority of decision tree-based methods is explained by the feature heterogeneity for tabular data, where neural networks struggle to learn the irregular target functions compared to decision tree-based methods grinsztajn2022tree.
Recently, LLMs have exhibited remarkable capabilities in natural language understanding and reasoning brown2020language; achiam2023gpt, sparking growing interest in applying LLMs to structured tabular data tasks dinh2022lift; hegselmann2023tabllm; yan2024making; wen2024supervised; nam2024tabular; han2024large; yan2025small. Typically, most of the existing methods first serialize tabular samples into natural language descriptions, and then tune LLMs or directly infer via in-context learning on these serialized data. Despite the success of LLMs on text-based tasks, leveraging LLMs to empower tabular predictions remains a challenging task. Through an in-depth investigation of prior LLM-based approaches for tabular data, we identify two inherent characteristics hindering prediction:
(i) Data perspective: To bridge the modality gap between unstructured text and structured tabular data with heterogeneous features, most existing methods serialize tabular samples into text formats. For instance, TabLLM hegselmann2023tabllm converts a sample into listed feature descriptions such as “The [column name] is [feature value]”, and demonstrates that this template facilitates very-few-shot classification by leveraging the semantic priors of column names and values already encoded in LLMs. However, many real-world tabular datasets anonymize feature names using placeholder symbols for privacy wen2024supervised (e.g., finance), reducing the effectiveness of serialization methods that rely on semantic feature descriptors. In addition, LLMs tend to be less sensitive to numerical features fang2024large; yan2024making. The predefined templates with inserted values often yield text formats that make LLMs struggle to understand the inherent interactions among different features. Beyond this, directly exposing the serialized tabular samples containing raw feature values to LLMs raises significant privacy concerns, especially in sensitive domains such as healthcare and finance, where data security is crucial yan2024protecting; yao2024survey. These challenges make it difficult to design a universally suitable serialization template for tabular data.
Table 1: Most existing LLM-based tabular prediction methods require access to tabular samples and either train the LLMs or perform inference via in-context learning. Many of them are restricted to constrained settings such as few-shot or classification-only settings. The applicability of each scenario is determined on the experimental setups in the original papers. For instance, if a method primarily focuses on for few-shot learning, we mark the few-shot setting as applicable ( ); if it can also be extended to full-data training, we additionally mark the full-data setting as applicable ( ).
Methods No LLM access sample required No LLM training required Applied scenario
No pre-training No fine-tuning Full data Few shot Classification Regression
TabLLM (2023) hegselmann2023tabllm
LIFT (2022) dinh2022lift
TP-BERTa (2024) yan2024making
GTL (2024) wen2024supervised
SERSAL (2025) yan2025small ✔(Unsupervised) ✔(Binary classification)
P2T (2024) nam2024tabular
FeatLLM (2024) han2024large
DeLTa (Ours)
(ii) Model perspective: One straightforward way to adapt LLMs to tabular prediction tasks is fine-tuning LLMs’ parameters. For example, LIFT dinh2022lift investigated the fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional tabular prediction methods. Despite the effectiveness, fine-tuning LLMs remains a challenging task even with recent parameter-efficient fine-tuning methods han2024parameter, especially in structured tabular data dinh2022lift. Another research line employs in-context learning by adding few-shot example demonstrations to the prompts without training, which is evidenced by LLMs’ impressive capacity to learn from demonstrations included as part of the prompt brown2020language. But most of these methods are inherently limited by the input length constraints of LLMs, restricting their use to few-shot, or classification-only tasks. Therefore, LLMs still struggle to make satisfactory predictions for tabular data. We summarize these related works in Table 1.
In this paper, we explore a novel direction of integrating LLMs into tabular data via logical decision tree rules loh2011classification as intermediaries. We propose DeLTa, a decision tree enhancer with LLM-derived rule for tabular prediction. Specifically, we first leverage the reasoning ability of LLMs to redesign an improved rule given a decision tree rule set, enforcing greater coherence among samples falling into the same leaf node. The newly generated rule is used to approximate a sample-specific error correction vector to calibrate the prediction of original decision trees in the direction of “errors” reducing. The proposed DeLTa could well solve the aforementioned challenges: Solving the data issue: Unlike serialization methods that convert each sample into unnatural text formats, decision tree rules are composed of simple comparisons between feature values and thresholds, forming logical, interpretable structures that can be naturally expressed in text without relying on semantic columns names. In addition, decision tree rules represent global feature space partitioning rule rather than individual samples, which helps mitigate privacy concerns by avoiding exposure of sample-level information. Solving the model issue: Moreover, the powerful reasoning ability of LLMs can be leveraged to redesign decision tree rules and help trees with aggregating their decisions, rather than directly using LLMs to generate label predictions. Notably, DeLTa avoids serialize tabular data into natural language format, and does not require additional domain-specific expertise or semantic information, such as explicit feature names and detailed task background knowledge. Furthermore, DeLTa can be applied in full data learning setting without LLM fine-tuning.
The contributions of this paper include: 1) We investigated the prior LLM-based methods for tabular prediction and explore a novel direction of integrating LLMs into tabular data via refining the decision tree rules, without directly accessing to the data itself, addressing the inherent issue w.r.t. data perspective and model perspective. 2) We propose a decision tree enhancer with LLM-derived rule for tabular prediction, DeLTa, which utilizes LLMs to refine a set of decision tree rules derived on decision trees trained on multiple train subsets, where the newly generated rule is used to infer the sample-wise error correction vector to calibrate the output of original decision trees. 3) We conduct extensive experiments on various tabular benchmarks and competing benchmark algorithms, and comprehensive results along with analysis and visualizations demonstrate our effectiveness.
2 Related work
Machine learning for tabular prediction. The development of effective algorithms for predictive modeling on tabular data has been a longstanding research topic. In the early days, decision tree-based methods (e.g. XGBoost chen2016xgboost, CatBoost prokhorenkova2018catboost) were found to be particularly suitable for tabular data. More recently, inspired by the success of of deep learning in computer vision (CV) he2016deep and natural language processing (NLP) devlin2018bert, numerous methods have been proposed for tabular data to accomplish tabular prediction tasks. These works mainly include MLP-like models wang2021dcn; gorishniy2021revisiting; klambauer2017self, attention-based architectures song2019autoint; gorishniy2021revisiting; somepalli2021saint; arik2021tabnet; huang2020tabtransformer, and retrieval-augmented architectures ye2025revisiting; somepalli2021saint; kossen2021self; gorishniy2023tabr. Among these works, DCN V2 wang2021dcn is an architecture that consists of an MLP-like module and a feature crossing module; AutoInt song2019autoint leveraged the Transformer architecture to capture inter-column correlations; FT-Transformer gorishniy2021revisiting further enhanced AutoInt’s performance through improved token embeddings; ModernNCA ye2025revisiting makes predictions based on the relationships with neighbors in a learned embedding space. Recently, another line of research has tried to use additional information outside target dataset to enhance tabular data prediction. XTab zhu2023xtab pretrains Transformer on a variety of datasets for cross-table pretraining. TabPFN hollmann2025accurate; hollmann2023tabpfn, which is pretrained on a large set of synthetic datasets, serves as a foundation model for small to medium-sized tabular data.
Large language models for tabular prediction. Motivated by the impressive success of LLMs brown2020language; achiam2023gpt; guo2024ds; guo2025optimizing, another promising study has attempted to harness the rich prior knowledge encapsulated by LLMs in tabular prediction tasks. LIFT dinh2022lift investigated the performance of fine-tuned GPT-3 models brown2020language on tabular data, revealing that the performance of fine-tuned LLMs was roughly on par with traditional solutions. Extending this line of research, TabLLM hegselmann2023tabllm employed T0 sanh2022multitask as the base LLM and demonstrated competitive performance of fine-tuned LLMs in very few-shot scenarios. Additionally, recent efforts have focused on pre-training LLMs over diverse tabular datasets from different domains. Among these, TP-BERTa yan2024making introduces a tabular-specific tokenization scheme, enabling a single pre-trained language model to generalize across multiple tabular datasets after further fine-tuning. GTL wen2024supervised further promotes comprehensive instruction-following capabilities for both zero-shot and in-context learning with a limited number of examples. Despite their promising results, these approaches typically require training or fine-tuning LLMs. Instead of directly fine-tuning LLMs, P2T nam2024tabular leverages unlabeled data correlated with target data expressed in natural language form and prompts LLMs for few-shot semi-supervised learning. Recently, Summary manikandan2023language proposed a boosting framework that treats LLMs as weak learners for tabular prediction, particularly for tasks involving small numbers of data points. FeatLLM han2024large prompts LLMs with serialized training examples to generate new features for few-shot classification tasks. SERSAL yan2025small explores to use synergy learning with FT-Transformer to enhance the noisy annotations generated by LLMs for binary classification tasks in unsupervised manner. This method also needs to fine-tune the LLMs. Going beyond tabular prediction, Nam et al. nam2024optimized leverages LLMs to optimize the feature generator with tree rules in the field of feature engineering zhang2023openfe, which we do not consider as a close related work to ours as the primary goal, motivation, and methodology are different.
Although prior works demonstrate that LLMs can make predictions for tabular tasks, the majority of these approaches rely on first converting tabular data into natural language descriptions, i.e., serialization, which often produces unnatural texts that differ from how humans might describe the data manikandan2023language. And it is also challenging to design such suitable template. In addition, these methods suffer from two limitations that either need to train LLMs or are restricted to constrained settings such as few-shot or classification-only settings. Table 1 summarizes these related works. The above limitations raise a fundamental question: can we harness the capabilities of LLMs without fine-tuning and without serializing tabular data, to improve full-data tabular prediction tasks across both classification and regression?
3 Preliminaries
Problem formulation. Denote a tabular dataset $\mathcal{D}\!=\!\{(x_{i},y_{i})\}_{i=1}^{N}$ as a collection of $N$ samples, where $(x_{i},y_{i})$ is the $i$-th data pair with $x_{i}\in\mathbb{X}$ representing input features and $y_{i}\in\mathbb{Y}$ the corresponding label. Concretely, we have $x_{i}\!=\!(x_{i}^{(num)},x_{i}^{(cat)})$ where $x_{i}^{(num)}$ and $x_{i}^{(cat)}$ represent the numerical and categorical features respectively. We consider supervised tabular prediction tasks: binary classification $\mathbb{Y}\!=\!\{0,1\}$, multiclass classification $\mathbb{Y}\!=\!\{1,...,c\}$ and regression $\mathbb{Y}\!=\!\mathbb{R}$. For data splits, $\mathcal{D}_{train}$ denotes training set for model training, $\mathcal{D}_{val}$ validation set for early stopping and hyperparameter tuning, and $\mathcal{D}_{test}$ test set for final evaluation. The goal is to obtain an accurate model $G:\mathbb{X}\to\mathbb{Y}$ trained on $\mathcal{D}_{train}$, that minimizes the expected loss $\mathbb{E}[\mathcal{L}(G(x),y)]$. Here, $\mathcal{L}$ is a smooth loss function (e.g., mean squared error or cross-entropy) and $G$ can be any tabular predictive model.
Decision trees (DTs). A decision tree (DT) loh2011classification is a hierarchical model that recursively partitions the input feature space into disjoint regions, i.e., leaf nodes, through a series of axis-aligned splits. For a given partitioning rule $r$, the prediction function $f$ can be formally expressed as:
$f(x|\mathcal{D}_{train},r)=\sum_{l=1}^{\text{Node}(r)}\lambda_{l}\cdot\mathbb{I}(x\in L_{l}(r)),$ (1)
where $\text{Node}(r)$ denotes the number of leaf nodes in the tree, $L_{l}(r)$ denotes the $l$-th leaf node containing a set of $x$ from $\mathcal{D}_{train}$, $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r)$ based on rule $r$, and $\lambda_{l}:\mathbb{X}\to\mathbb{Y}$ represents leaf-specific prediction values. Typically, $\lambda_{l}$ will be the average of corresponding labels for the leaf node in regression tasks, or the empirical class distribution in classification tasks. The structure of the tree is governed by a feature index vector $\ell$ and a threshold vector $\tau$, which together define the decision path containing a series of internal nodes from the root to each leaf node. Specifically, decision tree is constructed by recursively applying binary tests of the form $x^{\ell_{j}}\leq\tau_{j}$ at internal nodes, where $j$ indexes internal node along the root-to-leaf path, $\ell_{j}$ specifies feature index for $j$-th internal node split, $x^{\ell_{j}}$ denotes the corresponding feature value, and $\tau_{j}$ defines corresponding splitting threshold. In this work, we adopt CART loh2011classification as our decision tree implementation due to two primary reasons: (i) it offers high interpretability and can be expressed using a simple if-else syntax, (ii) tree-based models, often ensembles of simple decision trees like CART, outperform deep learning approaches in numerous tabular prediction tasks grinsztajn2022tree; gorishniy2021revisiting.
4 Proposed method: DeLTa
We propose a novel approach, Decision Tree Enhancer with LLM-derived Rule for Tabular Prediction (DeLTa), understanding the underlying rule logic encoded by decision trees and generating improved rules for tabular prediction via LLM, without requiring any fine-tuning of the LLM. This strategy introduces a fundamentally distinct interface for extending LLM capabilities to tabular tasks, departing from the majority of conventional data-to-text serialization pipelines that enforce LLMs to understand the serialized tabular data. In the following, we will elaborate on LLM-based decision tree rules refinement in Section 4.1; then, we will provide the calibration method for original decision trees via new generated rule by LLM in Section 4.2; we also give the overall implementation of ours in Section 4.3. An overview of the proposed framework is depicted in Fig. 1.
4.1 LLM-based decision tree rules refinement
The decision tree rules refinement procedure consists of the following stages: (i) decision tree rules initialization, that constructs a diverse decision tree rule set, and (ii) rule understanding via LLM, that leverages the LLM to understand and refine the original rule set to obtain a new rule.
Decision tree rules initialization. To mitigate the risk of overfitting for a single decision tree expert ho1998random, the classical Random Forest breiman2001random algorithm trains diverse decision tree experts $\{f_{k}\}_{k=1}^{K}$ on different subsets of $\mathcal{D}_{train}$, and uses the ensemble of $\{f_{k}\}_{k=1}^{K}$ as prediction:
$F(x)=\frac{1}{K}\sum_{k=1}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k}),$ (2)
where $F(x)$ denotes the label prediction produced by the Random Forest, $K$ is the number of experts, $\mathcal{D}_{train}^{k}$ denotes the subset, $r_{k}$ denotes the rule for expert $f_{k}$ derived from Eq. 1, and $\mathcal{R}=\{r_{k}\}_{k=1}^{K}$ denotes a source decision tree rule set. In terms of decision tree rule, each rule represents a logical comparison between features and thresholds displayed in semantically rich syntax, which thus can be used to partition the feature space into disjoint regions (i.e., leaf nodes) with high interpretability. A well understanding of these rules could generate a better feature space partitioning method, thereby promoting statistical coherence among samples assigned to the same leaf node. Although Random Forest ensembles the outputs from all rules, the inherent relationships and interactions among rules in $\mathcal{R}$ are ignored. With the development of $K$, analyzing these independent rules is gradually becoming more and more difficult, let alone utilizing these rules to partition feature space. To this end, we propose to leverage LLMs to analyze and summarize the rule set $\mathcal{R}$ into a refined rule due to the powerful logical reasoning ability of LLM.
Figure 1: The DeLTa framework. As shown in the main objective, we calibrate the output of original decision tree experts $F(x)$ in the direction of “errors” reducing. Subfig (a) describes the process of refining decision tree rules with LLM, and subfig (b) details the refined rule-guided error correction for decision trees.
Rule understanding via LLM. Our objective is to use an LLM to generate a new rule $r^{*}$ given original rule set $\mathcal{R}$. Specifically, we construct a prompt $p$ for the LLM that guides it to understand the diverse decision tree rule set $\mathcal{R}$ and synthesize the redesigned rule from the aggregated diverse knowledge. To effectively leverage the reasoning ability of LLM, the prompt $p$ (please see Appendix A.4) is designed to include: meta information $p_{meta}$ that describes the task objective, the extracted decision tree rule description $p_{rule}$ that contains $\mathcal{R}$, and the requirement $p_{requirement}$ for rule refinement. The formulation of generating $r^{*}$ is given by:
$r^{*}=LLM(p)=LLM(p_{meta}\oplus p_{rule}\oplus p_{requirement}),$ (3)
where $\oplus$ is the concatenation of each prompt, and this process is achieved through querying the LLM, without requiring any fine-tuning. Notably, our approach does not require additional domain-specific expertise or semantic information, such as explicit feature names or detailed task background knowledge. This makes it applicable to scenarios where such information is unavailable or incomplete. Instead of relying on the intrinsic domain knowledge of tabular tasks, such as those in healthcare or finance, we frame the problem purely as a domain knowledge-agnostic machine learning task. Our method leverages LLMs to reason decision tree rule set $\mathcal{R}$ into an rule $r^{*}$, rather than directly inferring from serialized tabular data. To verify that LLM could generate a better rule, where samples grouped within the same leaf node exhibit greater statistical similarity, we compute the intra-node sample distance over all leaf nodes partitioned by $r^{*}$, and observe that the distance of $r^{*}$ is lower than original rule $r\in\mathcal{R}$, as illustrated in Fig. 2. Next, we will elaborate on how to leverage the new rule $r^{*}$ generated by LLM to enhance tabular predictions. [Uncaptioned image] Figure 2: Average intra-node distance comparison.
4.2 Refined rule-guided decision tree error correction
The refined rule $r^{*}$ is summarized over the rule set $\mathcal{R}$, and thus provides a potentially better strategy that partitions the feature space of $\mathcal{D}_{train}$ into disjoint leaf nodes $\{L_{l}(r^{*})\}_{l=1}^{\text{Node}(r^{*})}$, with each leaf node containing a subset of $x$ from $\mathcal{D}_{train}$ and $\text{Node}(r^{*})$ denoting the number of leaf nodes. To use $r^{*}$ for making predictions, the simplest strategies include: (1) treating it as a standalone decision tree by assigning the input $x$ to a specific leaf node, or (2) appending it to the existing ensemble of decision trees $F(x)$. However, the former may be too isolated, as it does not leverage the predictive power of the existing ensemble, while the latter may be too diluted, as it does not fully utilize the refined guidance provided by LLMs. This work introduces a novel way to leverage the power of $r^{*}$ together with $F(x)$ inspired by the intuition that estimating the difference between predictions and labels, i.e., the residual errors is often easier than directly predicting the ground-truth labels zhang2005boosting; he2016deep. Intuitively, we first use $r^{*}$ to partition data samples into its leaf nodes, expecting the samples in the same node are similar to each other. Motivated by the famous Gradient Boosting algorithm friedman2001greedy; friedman2002stochastic, for the samples in each node, we learn a specific mapping function that predicts residual errors of $F(x)$ for the samples in that node. These learned functions are subsequently used to adjust the predictions made by $F(x)$, leading to improved accuracy. Since $r^{*}$ is designed to group similar samples effectively, the residuals within each leaf node are expected to be more structured and thus easier to predict.
Formally, we introduce the definition of the direction of prediction “errors” increasing and reducing, and the definition of a new set derived from $\mathcal{D}_{train}$ used for fitting the mapping functions as follows.
Definition 1. Prediction “errors” increasing and reducing. For a given training sample $x$, we approximate the "error" using the gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$, where $\mathcal{L}$ is the loss function discussed in Section 3. Intuitively, shifting $F(x)$ in the direction of this gradient (“Errors” Increasing) would likely worsen the model’s performance, analogous to how traditional gradients indicate a direction in parameter space that increases the loss ruder2016overview. Therefore, it is natural that we should calibrate $F(x)$ in the direction of negative gradient $\nabla_{F(x)}\mathcal{L}(F(x),y)$ (“Errors” Reducing).
Definition 2. Gradient set $\mathcal{D}_{train}^{\nabla}$. $\mathcal{D}^{\nabla}_{train}=\{(x,-\nabla_{F(x)}\mathcal{L}(F(x),y))|(x,y)\in\mathcal{D}_{train}\}$ is derived from train set $\mathcal{D}_{train}$ and stores training sample features and their corresponding negative gradients of the loss function with respect to the output of $F(x)$. The only difference between $\mathcal{D}_{train}$ and $\mathcal{D}_{train}^{\nabla}$ lies in the labels, which are replaced by the negative gradients in $\mathcal{D}_{train}^{\nabla}$, while $r^{*}$ induces the same leaf node partitioning of the feature space for both datasets.
After obtaining the Gradient Set $\mathcal{D}_{train}^{\nabla}$, we apply a learnable Gradient Net $\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ to model the mapping between $x$ and the corresponding negative gradient within $\mathcal{D}_{train}^{\nabla}$, where each leaf node $L_{l}(r^{*})$ corresponds to one specific mapping function $\phi_{l}(\cdot;\theta_{l}):x\to-\nabla_{F(x)}\mathcal{L}(F(x),y)$. The training process for $\phi_{l}$ is as follow:
$\min_{\theta_{l}}\sum_{x\in L_{l}(r^{*})}\|\phi_{l}(x;\theta_{l})-(-\nabla_{F(x)}\mathcal{L}(F(x),y))\|_{2}^{2},$ (4)
where $\phi_{l}$ could be easily implemented by conventional machine learning model, such as CART loh2011classification, and $\phi_{l}$ over different leaf nodes are trained separately.
At inference time, given $x$ from $\mathcal{D}_{test}$, the prediction of Gradient Net $\phi$ could be produced as a sample-specific error correction vector $\Delta_{x}$. The final prediction $F^{*}(x)$ could be obtained by adding $\Delta_{x}$ to $F(x)$, that moves $F(x)$ in the direction of negative gradient (“Errors” Reducing) by one step:
$\Delta_{x}=\eta*\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})=\eta*\sum_{l=1}^{\text{Node}(r^{*})}\phi_{l}(x;\theta_{l})\cdot\mathbb{I}(x\in L_{l}(r^{*})),$ (5)
$F^{*}(x)$ $=F(x)+\Delta_{x},$ (6)
where $\mathbb{I}(\cdot)$ is an indicator function determining if $x$ of interest would be assigned to leaf node $L_{l}(r^{*})$ based on rule $r^{*}$, $\eta\in\mathbb{R}_{+}$ is the hyperparameter similar to learning rate controlling the step size, and $\Delta_{x}$ approximates $\eta*\nabla_{F(x)}\mathcal{L}(F(x),y)$. We now explain the proposed error correction strategy by giving a closer look at Eq. 6.
Proposition 1
Let $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ denote the expected loss, where $F(x)=\frac{1}{K}\sum_{k}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k})$ and each $f_{k}$ corresponds to a decision tree rule $r_{k}$ from the expert-derived rule set $\mathcal{R}=\{r_{k}\}_{k=1}^{K}$. Given a prompt $p$ that contains $\mathcal{R}$, the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ could be decreased by querying the LLM with $p$ to generate a refined rule $r^{*}$, which enables approximation of a sample-specific error correction vector $\Delta_{x}$ to guide prediction $F(x)$ in the direction of “errors” reducing.
The proof of Proposition 1 is provided in Appendix A.5. To reduce the expected loss, we can approximate and apply negative gradient-based error correction vector $\Delta_{x}$ to $F(x)$.
4.3 Framework overview and discussion
Framework overview. In terms of training stage, the procedure begins by training a set of different decision tree experts (using CART for each expert) and obtaining the ensemble predictions based on Eq. 2. We then extract the corresponding decision tree rule set $\mathcal{R}$. The set of rules is used to construct a prompt $p$ for LLM, which returns an improved rule $r^{*}$ (Eq. 3). Subsequently, we compute the negative gradient for each training sample to construct the Gradient Set $\mathcal{D}_{train}^{\nabla}$, and then train the Gradient Net $\phi$ based on $\mathcal{D}_{train}^{\nabla}$ and $r^{*}$ (Eq. 4). During the test stage, we can achieve the preliminary prediction result $F(x)$ for each test sample $x$ based on the random forest; and also compute the sample-specific error correction vector $\Delta_{x}$ by feeding the test sample to Gradient Net. Now the final prediction can be corrected using Eq. 6. The complete training and inference pipeline is summarized in Algorithm 1 of Appendix A.3.
Discussion. To sum up, fed with the expert-derived rules $\mathcal{R}$, LLM is prompted to summarize multiple rules and redesign a new rule $r^{*}$, which thus considers the inherent relationships and interactions among multiple independent rules. Benefiting from the rule-based prompt, we neither need to fine-tune LLM nor expose the serialized samples into LLM. Due to the powerful reasoning ability of LLM, the refined rule $r^{*}$ that summarizes multiple rules is more effective and also convenient for us to use. For example, we can not only use the rule to make decision but also to partition the feature space of training set. The latter means that we can build a set of disjoint leaf nodes, where samples grouped within the same leaf node exhibit greater statistical similarity and thus can be used to train network. Here, we use the samples in the same leaf node to model the mapping between sample and the corresponding negative gradient, which is usually easier than directly fitting the relationship between samples and the labels. Our proposed DeLTa enables a principled integration of expert knowledge and LLM-based rule synthesis for enhanced performance, which paves a new way for LLM-based tabular data learning.
5 Experiments
Datasets. We consider a variety of supervised tabular prediction tasks with heterogeneous features, including binary classification, multiclass classification, and regression. Specifically, the tabular datasets include: Blood (BL) blood_transfusion_service_center_176, Credit (CR) statlog_(german_credit_data)_144, Car bohanec1988knowledge, Bank (BA) Moro2014ADA, Adult (AD) kohavi1996scaling, Jannis (JA) grinsztajn2022tree, Cpu_act (CP) grinsztajn2022tree, Credit_reg (CRR) grinsztajn2022tree, California_housing (CA) pace1997sparse, House_16H (HO) grinsztajn2022tree, Fried (FR) breiman1996bagging, Diomand (DI) chen2023trompt. The dataset properties and data pre-processing details are summarized in Appendix A.1. Following previous studies gorishniy2021revisiting, we use Accuracy (higher is better) to evaluate binary and multiclass classification tasks, Normalized Root Mean Squared Error (NRMSE) (lower is better) to evaluate the regression tasks.
Baselines and implementation details. We conduct a comparative analysis between DeLTa and other prominent methods in the field of tabular prediction. Specifically, we include LLM-based methods such as TabLLM hegselmann2023tabllm, LIFT dinh2022lift, TP-BERTa yan2024making, GTL wen2024supervised, P2T nam2024tabular, FeatLLM han2024large. For LIFT, we employ two versions, i.e., fine-tuning LLMs (LIFT) and in-context learning (LIFT-ICL) versions, as provided in their original paper. In addition, we compare DeLTa with non LLM-based methods, including KNN cover1967nearest, CART loh2011classification, MLP lecun2015deep, Random Forest breiman2001random, XGBoost chen2016xgboost, CatBoost prokhorenkova2018catboost, FT-Transformer gorishniy2021revisiting, TabPFN hollmann2025accurate, ModernNCA ye2025revisiting. More details about the baseline models can be found in Appendix A.2. Implementation details of DeLTa including hyper-parameters are provided in Appendix A.3. For LLM usage, DeLTa adopts GPT-4o as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix A.6 for more results). To prevent LLMs from generating task-irrelevant output, we query LLM 10 times by default and use the average. In replicating baselines, GPT-4o is utilized for in-context learning. For baselines require LLM fine-tuning, we employ the LLMs used in their original paper: TabLLM uses T0 sanh2022multitask; we use the API provided by OpenAI to perform black-box GPT-3.5 fine-tuning for LIFT, where the fine-tuning method is suggested by the original paper; TP-BERTa employs RoBERTa liu2019roberta; GTL uses the 13B version of LLaMA 2 touvron2023llama. The experiments are run on NVIDIA A100-PCIE-40GB GPU.
Table 2: Test Accuracy ($\uparrow$) performance of DeLTa and LLM-based baseline methods on classification tasks. “# Num” indicates the number of training samples. “Tab” indicates whether the LLM requires access to tabular samples. “Extra” indicates whether the method needs to use extra data samples outside the training set. “Feat” indicates whether the feature names are needed. “Tune” indicates whether LLM fine-tuning is needed.
# Num Method (LLM used) Additional requirements Datasets
Tab Extra Feat Tune BL $\uparrow$ CR $\uparrow$ Car $\uparrow$ BA $\uparrow$ AD $\uparrow$ JA $\uparrow$ Average $\uparrow$
All TabLLM (T0) 0.761 0.737 0.845 0.906 0.864 0.691 0.801
All LIFT (GPT-3.5) 0.689 0.691 0.729 0.825 0.81 0.647 0.732
All TP-BERTa (RoBERTa) 0.761 0.730 0.826 0.916 0.844 0.659 0.789
All GTL (LLaMA2) N/A N/A N/A N/A N/A N/A N/A
All LIFT-ICL (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
All P2T (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
All FeatLLM (GPT-4o) 0.768 0.701 0.589 0.887 0.842 0.540 0.721
All DeLTa (GPT-4o) 0.829 0.783 0.836 0.908 0.868 0.705 0.822
256 TabLLM (T0) 0.344 0.690 0.471 0.785 0.787 0.373 0.575
256 LIFT (GPT-3.5) 0.651 0.682 0.431 0.724 0.773 0.398 0.610
256 TP-BERTa (RoBERTa) 0.721 0.677 0.408 0.853 0.772 0.464 0.649
256 GTL (LLaMA2) N/A N/A N/A N/A N/A N/A N/A
256 LIFT-ICL (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
256 P2T (GPT-4o) N/A N/A N/A N/A N/A N/A N/A
256 FeatLLM (GPT-4o) 0.729 0.658 0.459 0.740 0.772 0.473 0.639
256 DeLTa (GPT-4o) 0.732 0.713 0.736 0.767 0.778 0.476 0.700
64 TabLLM (T0) 0.245 0.673 0.386 0.602 0.785 0.268 0.493
64 LIFT (GPT-3.5) 0.643 0.643 0.500 0.728 0.748 0.430 0.615
64 TP-BERTa (RoBERTa) 0.648 0.646 0.419 0.840 0.761 0.414 0.621
64 GTL (LLaMA2) 0.613 0.638 0.512 0.719 0.750 0.414 0.608
64 LIFT-ICL (GPT-4o) 0.348 0.538 0.326 0.469 0.613 0.251 0.424
64 P2T (GPT-4o) 0.368 0.542 0.334 0.569 0.634 0.243 0.448
64 FeatLLM (GPT-4o) 0.697 0.643 0.508 0.723 0.749 0.450 0.628
64 DeLTa (GPT-4o) 0.732 0.663 0.615 0.733 0.758 0.457 0.660
Figure 3: Test NRMSE ($\downarrow$) performance of DeLTa and LLM-based baseline methods on regression tasks.
5.1 Main Results
Comparison with LLM-based methods. We evaluated the performance of DeLTa against mainstream LLM-based baseline methods on tabular classification tasks, as illustrated in Table 2. In addition to standard prediction tasks under full-data regimes, we also provide results under low-data regimes (few-shot learning), as some of these LLM-based baseline methods are constrained by the limited input length of LLMs and are thus restricted to low-data settings. Following prior work han2024large, we simulate the few-shot setting by randomly sampling a fixed number of training examples, where the number of shots denotes the total number of selected samples. Considering some of the LLM-based baseline methods are applicable to regression tasks, we further evaluate DeLTa on regression under both full-data and low-data regimes settings, as illustrated in Fig. 3. Overall, DeLTa achieves the highest average performance in all settings, including classification and regression, in both full and low-data regimes. While DeLTa is primarily designed for full-data tabular prediction, our results demonstrate that it also performs competitively in low-data scenarios. Furthermore, compared to existing LLM-based tabular prediction methods, the proposed DeLTa requires no additional requirements for LLMs as illustrated in Table 2. It also aligns with our claims that we have solved the two inherent issues for prior LLM-based approaches: (i) the proposed DeLTa avoids reliance on column names and tabular data serialization, and thus will protect sample-level privacy information, and (ii) rather than directly using LLMs to generate label predictions, DeLTa utilizes the powerful reasoning ability of LLMs to enhance decision trees, enabling effective tabular prediction in both full-data and low-data regimes without requiring LLM fine-tuning.
Computational efficiency. We compare the computational cost between ours and LLM-based methods in Table 3. The results show that our proposed DeLTa is computationally efficient in both training and inference stages. This efficiency stems from key properties: (i) DeLTa utilizes the reason ability of LLM without fine-tuning LLM, (ii) DeLTa avoids querying LLMs to generate predictions for individual samples. Instead, it only needs to query LLM via API to generate one refined decision tree rule for one dataset, enabling significantly more efficient use of LLM resources.
Table 3: Runtime in seconds of DeLTa and other LLM-based baseline methods for the training and inference phase, conducted on Adult dataset. “# Num” indicates number of training samples. †These methods require querying LLMs for each test sample. * These methods do not require querying LLMs at inference phase.
# Num Stage TabLLM† LIFT† TP-BERTa† GTL† LIFT-ICL† P2T† FeatLLM* DeLTa* (Ours)
All Train 177371.37 153689.14 1319.35 N/A N/A N/A 1231.22 35.20
Inference 179.21 90149.48 1.58 N/A N/A N/A 0.14 0.09
64 Train 288.61 191.75 316.40 N/A N/A N/A 1047.55 23.09
Inference 170.42 89684.76 5.01 153501.02 176726.13 200258.24 0.14 0.08
Comparison with conventional methods. We further compare the performance of DeLTa and other conventional tabular models on full data regimes averaged over all datasets, as illustrated in Fig. 4. Full results are provided in Appendix A.8. The results show that DeLTa also performs well compared with conventional tabular models on both classification and regression tasks.
Figure 4: Test performance of DeLTa and non-LLM baseline methods on classification and regression tasks. Here, “RF” denotes Random Forest, “FT-T” denotes FT-Transformer, “MNCA” denotes ModernNCA.
5.2 Further analysis
Ablation study. We further conduct ablation study to demonstrate the effectiveness of key components of DeLTa, as illustrated in Table 4, by comparing DeLTa against three variants: (i) DeLTa w/o “RR”, that replaces the LLM generated $r^{*}$ used in Gradient Net $\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ with $r\in\mathcal{R}$ derived from Random Forest $F(x)$, (ii) DeLTa w/o “EC”, that directly obtains the label prediction for test $x$ by utilizing $f(x|\mathcal{D}_{train},r^{*})$ (Eq. 1) with the refined rule $r^{*}$, (iii) Random Forest, which makes decision with $F(x)$ in Eq. 2. The results show that the decision tree rule refinement via LLM is important, and the error correction strategy could further enhance tabular prediction performance.
Table 4: Ablation study on the different components in DeLTa; see full results in Appendix A.10. “RR” denotes the rule refinement process, and “EC” denotes the error correction process.
Variants CR $\uparrow$ CA $\downarrow$
DeLTa (Ours) 0.7829 0.3507
w/o “RR” 0.7257 0.5443
w/o “EC” 0.7657 0.4016
Random Forest 0.7371 0.5057
Additional results. Fig. 5 shows the label predictions of ours: $F(x)+\Delta_{x}$ and ours w/o $\Delta_{x}$: $F(x)$. We can observe that two categories of samples are overlapped in some regions. $F(x)$ struggles to distinguish them, but DeLTa could correct the prediction errors of $F(x)$ to enhance the prediction for such complex data patterns.
Figure 5: Visualization of label prediction of DeLTa w/ and w/o error correction vector $\Delta_{x}$ on BA dataset.
We incorporate the sensitivity analysis for hyperparameters in Appendix A.11. Additional insights and analysis regarding the LLM-refined rule can be found in Appendix A.12.
6 Conclusion
In this paper, we propose DeLTa, a decision tree enhancer with LLM-derived rule for tabular prediction, to solve the two challenges w.r.t. data perspective and model perspective in prior LLM-based approaches for tabular prediction. DeLTa avoids reliance on feature names and tabular data serialization, and it utilizes the powerful reasoning ability of LLMs to enhance decision trees, enabling effective tabular prediction in both full-data and low-data regimes without requiring LLM fine-tuning. The empirical results demonstrated the effectiveness of DeLTa for tabular prediction tasks. Our work can shed some light on developing better algorithms for similar tasks.
Acknowledgments and Disclosure of Funding
The authors would like to thank the anonymous referees for their valuable comments. In this work, Hangting Ye, Jinmeng Li, Dandan Guo and Yi Chang are supported by the National Key R&D Program of China under Grant (No. 2023YFF0905400) and the National Natural Science Foundation of China (No. U2341229, No. 623B2043, No. 62306125).
Part of this work was carried out during Dandan Guo’s visit to King Abdullah University of Science and Technology (KAUST).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the claims made in the abstract and introduction accurately reflect the paper’s contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, we discussed the limitations of the work in Appendix A.14.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, we provide the full set of assumptions and a complete (and correct) proof in Appendix A.5.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, we provide the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Yes, we provide the access to the data and code with instructions to reproduce the main experimental results at https://github.com/HangtingYe/DeLTa.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, we specify all the training and test details necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report standard deviation in Appendix A.7.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, we provide information on the computer resources needed to reproduce the experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, the research conducted in the paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, we discuss the both potential positive societal impacts and negative societal impacts of the work performed in Appendix A.16.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, the creators or original owners of assets used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes, new assets introduced in the paper well documented and is the documentation provided alongside the assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Yes, we describe the usage of LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
A.1 Datasets details
We consider a variety of supervised tabular prediction tasks with heterogeneous features, including classification and regression. Specifically, the tabular datasets include: Blood (BL), Credit (CR), Car, Bank (BA), Adult (AD), Jannis (JA), Cpu_act (CP), Credit_reg (CRR), California_housing (CA), House_16H (HO), Fried (FR), Diomand (DI). The dataset properties are summarized in Table 5. Our primary objective was to evaluate DeLTa’s effectiveness in the context of LLM-based tabular prediction, and to ensure fair comparison with existing methods in this category. Accordingly, the majority of our classification datasets were selected based on their widespread use in recent LLM-based tabular learning studies such as TabLLM and FeatLLM, both of which primarily focus on classification tasks. To extend our evaluation beyond classification, we also incorporated regression datasets from the TALENT benchmark [ye2024closerlookdeeplearning, liu2024talenttabularanalyticslearning], which has recently gained attention for assessing tabular regression performance. To ensure fair comparisons, we adhere to identical preprocessing procedures following TALENT benchmark [liu2024talenttabularanalyticslearning] for each dataset.
Table 5: Tabular dataset properties. “#objects” indicates the number of samples in the dataset. “#num. features” indicates the number of numerical features, and “#cat. features” indicates the number of categorical features.
BL CR Car BA AD JA CP CRR CA HO FR DI
#objects 748 1000 1728 45211 48842 83733 8192 16714 20640 22784 40768 53940
#num. features 4 7 0 7 6 54 21 10 8 16 10 6
#cat. features 0 13 6 9 8 0 0 0 0 0 0 3
metric Acc. Acc. Acc. Acc. Acc. Acc. NRMSE NRMSE NRMSE NRMSE NRMSE NRMSE
#classes 2 2 4 2 2 4
A.2 Baseline models details
We include LLM-based methods such as TabLLM, LIFT, LIFT-ICL, TP-BERTa, GTL, P2T, FeatLLM: (1) TabLLM employs T0 as the base LLM and investigates the performance of fine-tuned T0 model in very few-shot scenarios; (2) LIFT investigates the performance of fine-tuned GPT models on tabular data; (3) LIFT-ICL is another version of LIFT provided in the original paper, that utilizes in-context learning without fine-tuning by providing a few training examples; (4) TP-BERTa introduces a tabular-specific tokenization scheme, pretraining a single language model across multiple tabular datasets. Given target tabular datasets, the pretrained language model should be further fine-tuned; (5) GTL pretrains LLaMA across multiple tabular datasets to achieve zero-shot and in-context learning on target tabular dataset; (6) P2T leverages extra unlabeled samples outside the training set as demonstrations to conduct in-context learning; (7) FeatLLM prompts LLMs with a few training examples to generate new features to replace original features and then trains linear models for few-shot classification tasks. In replicating baselines, GPT-4o is utilized for in-context learning. For baselines require LLM fine-tuning, we employ the LLMs used in their original paper: TabLLM uses T0; we use the API provided by OpenAI to perform black-box GPT-3.5 fine-tuning for LIFT, where the fine-tuning method is suggested by the original paper; TP-BERTa employs RoBERTa; GTL uses the 13B version of LLaMA 2. The implementation of these methods is based on their official open-source code releases.
In addition, we compare DeLTa with non LLM-based methods, including KNN, CART, MLP, Random Forest, XGBoost, CatBoost, ResNet, FT-Transformer, Saint, TabPFN and ModernNCA. Among them, (1) KNN classifies or predicts a sample based on the majority vote or average of its $k$ nearest neighbors in the feature space; (2) CART is an implementation for the simplest decision trees; (3) MLP is a type of feedforward neural network composed of multiple layers; (4) Random Forest, (5) XGBoost and (6) CatBoost are the variants of decision tree-based methods; (7) ResNet is a deep neural network architecture that introduces skip connections to enable the training of very deep networks; (8) FT-Transformer is a token-based method which transforms features to embeddings and applies a series of attention-based transformations to the embeddings; (9) Saint is a token-based method that leverages row and column attention mechanisms for tabular data; (10) TabPFN generates a synthetic dataset with diverse distribution to pretrain a model, and then leverages the training samples of target dataset for inference. We use the version 2 of TabPFN since it can be applied to both classification and regression datasets; (11) ModernNCA makes predictions based on the relationships with neighbors in a learned embedding space. The implementation of these methods is based on the official open-source code releases from TALENT benchmark [liu2024talenttabularanalyticslearning].
A.3 DeLTa details
Algorithm 1 DeLTa training and inference workflow.
Input: Training set $\mathcal{D}_{train}$, test set $\mathcal{D}_{test}$ (without accessing to label), number of decision tree experts $K$, API of LLM;
Initialize source decision tree rule set $\mathcal{R}=[]$;
for $k=1$ to $K$ do // This loop can be executed in parallel.
  Train one decision tree expert $f_{k}$ on $\mathcal{D}_{train}^{k}$ with Eq. 1;
  Extract the corresponding rule $r_{k}$ of $f_{k}$;
  Update $\mathcal{R}\leftarrow\mathcal{R}\cup r_{k}$
end for
Obtain the ensemble prediction of decision tree experts $F(x)$ with Eq. 2;
Form prompt $p$ based on $\mathcal{R}$;
Query LLM with prompt $p$ and generate the new rule $r^{*}$ with Eq. 3;
Calculate Gradient Set $\mathcal{D}^{\nabla}_{train}=\{(x,-\nabla_{F(x)}\mathcal{L}(F(x),y))|(x,y)\in\mathcal{D}_{train}\}$;
Train Gradient Net $\phi$ based on Gradient Set with Eq. 4;
for $x$ in $\mathcal{D}_{test}$ do // Without accessing to label. And this loop can be executed in parallel.
  Estimate error correction vector $\Delta_{x}=\eta*\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ with Eq. 5;
  Computing final prediction $F^{*}(x)=F(x)+\Delta_{x}$ with Eq. 6;
end for
Return: Final prediction $\{F^{*}(x)|x\in\mathcal{D}_{test}\}$.
The DeLTa architecture remains consistent across all datasets. Specifically, the $F(x)$ is produced by the output of Random Forest. For LLM usage, DeLTa adopts GPT-4o via API as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Table 6 and Table 7 for more results). To prevent LLMs from generating task-irrelevant output, we query LLM 10 times by default and use the average. DeLTa needs to construct Gradient Set $\mathcal{D}^{\nabla}_{train}=\{(x,-\nabla_{F(x)}\mathcal{L}(F(x),y))|(x,y)\in\mathcal{D}_{train}\}$, which stores training sample features and their corresponding negative gradients of the loss function with respect to the output of $F(x)$. Here, $\mathcal{L}$ is a smooth loss function (e.g., mean squared error for regression or cross-entropy for classification). And $\nabla_{F(x)}\mathcal{L}(F(x),y)\in\mathbb{R}$ for regression, $\nabla_{F(x)}\mathcal{L}(F(x),y)\in\mathbb{R}$ for binary classification, and $\nabla_{F(x)}\mathcal{L}(F(x),y)\in\mathbb{R}^{c}$ for multiclass classification, where $c$ is the number of classes. The Gradient Net $\phi$ contains learnable leaf node-specific mapping function $\phi_{l}$, where $\phi_{l}$ is implemented by CART for classification and TabPFN for regression, and $\phi_{l}$ over different leaf nodes are trained separately. The number of leaf node for LLM generated rule $r^{*}$ will less than 30, and different $\phi_{l}$ can be trained in parallel, which will result in a low runtime cost as illustrated in Table 14. The training and inference pipeline is summarized in Algorithm 1.
A.4 Example prompt for refining decision tree rules
We leverage LLM to refine the decision tree rule set $\mathcal{R}$ to generate a better rule $r^{*}$, where samples grouped within the same leaf node exhibit greater statistical similarity. Therefore, we query LLM to generate a better rule to partition the feature space into disjoint regions (i.e., leaf nodes). Specifically, the prompt $p$ is designed to include: meta information $p_{meta}$ that describes the task objective, the extracted decision tree rule description $p_{rule}$ that contains $\mathcal{R}$, and the requirement $p_{requirement}$ for rule refinement. Please note that our approach does not require additional domain-specific expertise or semantic information, such as explicit feature names or detailed task background knowledge. This makes it applicable to scenarios where such information is unavailable or incomplete. Instead of relying on the intrinsic domain knowledge of tabular tasks, such as those in healthcare or finance, we frame the problem purely as an domain knowledge-agnostic machine learning task. Our method leverages LLMs to reason decision tree rule set $\mathcal{R}$ into an rule $r^{*}$, rather than directly inferring from serialized tabular data, and thus will protect sample-level privacy information. The prompt example is provided in Fig. 6.
You are an expert in tabular machine learning domain. I will provide the meta information, the CART tree rules about the prediction task. Please help me design a better rule for inference.
# Meta information about dataset.
    "name": "adult",
    "task_type": "binclass",
    "n_num_features": 6,
    "n_cat_features": 8,
    "train_size": 26048,
# CART tree rules (We provide the example of decision rule of one subtree from the Random Forest below.)
    Tree 1 rules:
    |--- feature_10 <= -0.59
    |   |--- feature_4 <= 4.31
    |   |   |--- feature_2 <= 0.94
    |   |   |   |--- feature_6 <= -1.82
    |   |   |   |   |--- class: 1.0
    |   |   |   |--- feature_6 >  -1.82
    |   |   |   |   |--- class: 0.0
    |   |   |--- feature_2 >  0.94
    |   |   |   |--- feature_3 <= 0.53
    |   |   |   |   |--- class: 1.0
    |   |   |   |--- feature_3 >  0.53
    |   |   |   |   |--- class: 1.0
    |   |--- feature_4 >  4.31
    |   |   |--- class: 1.0
    |--- feature_10 >  -0.59
    |   |--- feature_0 <= -0.74
    |   |   |--- class: 0.0
    |   |--- feature_0 >  -0.74
    |   |   |--- feature_3 <= 0.80
    |   |   |   |--- feature_2 <= 0.94
    |   |   |   |   |--- class: 0.0
    |   |   |   |--- feature_2 >  0.94
    |   |   |   |   |--- feature_10 <= 1.89
    |   |   |   |   |   |--- class: 0.0
    |   |   |   |   |--- feature_10 >  1.89
    |   |   |   |   |   |--- class: 1.0
    |   |   |--- feature_3 >  0.80
    |   |   |   |--- class: 1.0
    Tree 2 rules:
# CART tree rules end
Based on the above information, please learn the rules evolving process and help me design a better rule like what cart used for inference to achieve higher performance. Please not just copy, please refine these rules and create a new better one. The new rule aims to divide the training space into several regions, where each region is denoted by a unique leaf node id. The number of leaf nodes should no more than $x$. Please return the dict format of rule, the format should be strictly like:
self.tree = {
        "feature": 11,
        "threshold": -0.78,
        "operator": "<=",
        "left": {"id": "leaf_1"},
        "right": {
            "feature": 7,
            "threshold": -0.46,
            "operator": "<=",
            "left": {"id": "leaf_2"},
            "right": {"id": "leaf_3"},
Please note that each leaf id node can only appear once, for example, "id": "leaf_1" can only appear once. Thus you only need to return the leaf nodes, rather than the true predictions.
Figure 6: An example of the prompt template for the adult dataset. Here, $x$ denotes the maximum number of leaf nodes among all the provided decision trees.
A.5 Theoretical analysis
Proposition 2
Let $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ denote the expected loss, where $F(x)=\frac{1}{K}\sum_{k}^{K}f_{k}(x|\mathcal{D}_{train}^{k},r_{k})$ and each $f_{k}$ corresponds to a decision tree rule $r_{k}$ from the expert-derived rule set $\mathcal{R}=\{r_{k}\}_{k=1}^{K}$. Given a prompt $p$ that contains $\mathcal{R}$, the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ could be decreased by querying the LLM with $p$ to generate a refined rule $r^{*}$, which enables approximation of a sample-specific error correction vector $\Delta_{x}$ to guide prediction $F(x)$ in the direction of “errors” reducing.
Proof. The the expected loss $\mathbb{E}\left[\mathcal{L}(F(x),y)\right]$ after calibration is formulated as:
$\mathbb{E}\left[\mathcal{L}(F^{*}(x),y)\right]$ $=\mathbb{E}\left[\mathcal{L}(F(x)+\Delta_{x},y)\right]$ (7)
$\approx\mathbb{E}\left[\mathcal{L}(F(x),y)+\Delta_{x}*\nabla_{F(x)}\mathcal{L}(F(x),y)\right]$
$=\mathbb{E}\left[\mathcal{L}(F(x),y)+(-\eta*\nabla_{F(x)}\mathcal{L}(F(x),y))*\nabla_{F(x)}\mathcal{L}(F(x),y)\right]$
$=\mathbb{E}\left[\mathcal{L}(F(x),y)+(-\eta*\|\nabla_{F(x)}\mathcal{L}(F(x),y)\|_{2}^{2})\right]$
$<\mathbb{E}\left[\mathcal{L}(F(x),y)\right],$
where $\mathbb{E}\left[\mathcal{L}(F(x)+\Delta_{x},y)\right]\approx\mathbb{E}\left[\mathcal{L}(F(x),y)+\Delta_{x}*\nabla_{F(x)}\mathcal{L}(F(x),y)\right]$ is given by Taylor first-order expansion and $\mathcal{L}$ is a convex loss function (e.g., mean squared error or cross-entropy). Error correction vector $\Delta_{x}=\eta*\phi(x|\mathcal{D}^{\nabla}_{train},r^{*})$ approximates the $\eta*\nabla_{F(x)}\mathcal{L}(F(x),y)$ in the direction of “errors” reducing according to Eq. 4 in the original paper and $\eta\in\mathbb{R}_{+}$ is the hyperparameter. To reduce the expected loss, we can approximate and apply negative gradient-based error correction vector $\Delta_{x}$ to $F(x)$.
A.6 Varying LLM backbones
To investigate the impact of using different LLMs in our framework, given their distinct prior knowledge and reasoning abilities from being trained on various text corpora, we measured performance using not only GPT-4o but also the Qwen3-32B (open-source) as a backbone for comparison. The results in Table 6 and Table 7 indicate that while certain tasks show greater improvement with specific LLMs, the average performance is similar with different LLMs.
Table 6: Effects of various LLM backbones: GPT-4o vs. Qwen3-32B on classification tasks $\uparrow$ in full-data and low-data regimes. “# Num ” indicates the number of training samples.
[…]
All DeLTa (Qwen3-32B) 0.821 0.766 0.832 0.907 0.867 0.703 0.816
All DeLTa (GPT-4o) (Ours) 0.829 0.783 0.836 0.908 0.868 0.705 0.822
256 DeLTa (Qwen3-32B) 0.741 0.707 0.750 0.765 0.778 0.474 0.703
256 DeLTa (GPT-4o) (Ours) 0.732 0.713 0.736 0.767 0.778 0.476 0.700
128 DeLTa (Qwen3-32B) 0.744 0.687 0.679 0.749 0.778 0.456 0.682
128 DeLTa (GPT-4o) (Ours) 0.746 0.683 0.722 0.749 0.774 0.457 0.688
64 DeLTa (Qwen3-32B) 0.736 0.657 0.567 0.729 0.759 0.454 0.650
64 DeLTa (GPT-4o) (Ours) 0.732 0.663 0.615 0.733 0.758 0.457 0.660
Table 7: Effects of various LLM backbones: GPT-4o vs. Qwen3-32B on regression tasks $\downarrow$ in full-data and low-data regimes. “# Num ” indicates the number of training samples.
[…]
All DeLTa (Qwen3-32B) 0.114 0.788 0.367 0.541 0.201 0.129 0.357
All DeLTa (GPT-4o) (Ours) 0.116 0.780 0.351 0.529 0.200 0.130 0.351
256 DeLTa (Qwen3-32B) 0.203 0.854 0.606 0.811 0.358 0.293 0.521
256 DeLTa (GPT-4o) (Ours) 0.234 0.841 0.568 0.799 0.397 0.277 0.519
128 DeLTa (Qwen3-32B) 0.206 0.876 0.709 0.940 0.474 0.348 0.592
128 DeLTa (GPT-4o) (Ours) 0.247 0.857 0.641 0.855 0.512 0.319 0.572
64 DeLTa (Qwen3-32B) 0.289 0.920 0.737 0.944 0.598 0.406 0.649
64 DeLTa (GPT-4o) (Ours) 0.382 0.888 0.701 0.919 0.634 0.383 0.651
A.7 Full comparison results with LLM-based baseline methods
We provide the full results of DeLTa against mainstream LLM-based baseline methods in full-data and low-data regimes in Table 8 to Table 11. “# Num” indicates the number of training samples. “Tab” indicates whether the LLM requires access to tabular samples. “Extra” indicates whether the method needs to use extra data samples outside the training set. “Feat” indicates whether the feature names are needed. “Tune” indicates whether LLM fine-tuning is needed.
Table 8: Comparison with LLM-based methods averaged over all classification datasets $\uparrow$. Average standard deviations over all datasets are given.
Methods Additional requirements Classification, # Num
Tab Extra Feat Tune All 512 256 128 64 16
TabLLM (T0) 0.801.012 0.568.035 0.575.078 0.519.046 0.493.054 0.413.082
LIFT (GPT-3.5) 0.732.051 0.625.045 0.610.056 0.650.064 0.615.059 0.556.083
TP-BERTa (RoBERTa) 0.789.016 0.638.029 0.649.063 0.641.076 0.621.073 0.563.091
GTL (LLaMA2) N/A N/A N/A N/A 0.608.058 0.552.079
LIFT-ICL (GPT-4o) N/A N/A N/A N/A 0.424.107 0.382.081
P2T (GPT-4o) N/A N/A N/A N/A 0.448.060 0.304.072
FeatLLM (GPT-4o) 0.721.024 0.664.024 0.639.041 0.646.041 0.628.047 0.570.066
Ours (GPT-4o) 0.822.005 0.719.026 0.700.018 0.688.022 0.660.045 0.579.055
Table 9: Comparison with LLM-based methods averaged over all regression datasets $\downarrow$. Average standard deviations over all datasets are given.
Methods Additional requirements Regression, # Num
Tab Extra Feat Tune All 512 256 128 64 16
TabLLM (T0) N/A N/A N/A N/A N/A N/A
LIFT (GPT-3.5) 0.623.048 0.782.056 0.715.054 0.743.047 0.827.098 0.995.095
TP-BERTa (RoBERTa) 0.403.015 0.496.011 0.582.019 0.642.036 0.727.028 0.887.028
GTL (LLaMA2) N/A N/A N/A N/A 0.847.072 0.938.087
LIFT-ICL (GPT-4o) N/A N/A N/A N/A 0.820.084 1.025.084
P2T (GPT-4o) N/A N/A N/A N/A 0.935.062 1.042.064
FeatLLM (GPT-4o) N/A N/A N/A N/A N/A N/A
Ours (GPT-4o) 0.351.008 0.462.015 0.519.028 0.572.032 0.651.043 0.852.056
Table 10: Comparison with LLM-based methods on all classification datasets $\uparrow$. We also calculate the average relative improvement of DeLTa against the best baseline method ($\uparrow$ %). “# Num” indicates the number of training samples. This table serves as an extension of Table 8.
[…]
All TabLLM (T0) 0.761 0.737 0.845 0.906 0.864 0.691 0.801
All LIFT (GPT-3.5) 0.689 0.691 0.729 0.825 0.810 0.647 0.732
All TP-BERTa (RoBERTa) 0.761 0.730 0.826 0.916 0.844 0.659 0.789
All GTL (LLaMA2) - - - - - - -
All LIFT-ICL (GPT-4o) - - - - - - -
All P2T (GPT-4o) - - - - - - -
All FeatLLM (GPT-4o) 0.768 0.701 0.589 0.887 0.842 0.540 0.721
All DeLTa (GPT-4o) (Ours) 0.829 0.783 0.836 0.908 0.868 0.705 0.822 ($\uparrow$ 2.6%)
512 TabLLM (T0) 0.240 0.712 0.447 0.812 0.788 0.412 0.568
512 LIFT (GPT-3.5) 0.673 0.692 0.503 0.735 0.710 0.437 0.625
512 TP-BERTa (RoBERTa) 0.716 0.694 0.275 0.889 0.767 0.487 0.638
512 GTL (LLaMA2) - - - - - - -
512 LIFT-ICL (GPT-4o) - - - - - - -
512 P2T (GPT-4o) - - - - - - -
512 FeatLLM (GPT-4o) 0.736 0.681 0.544 0.745 0.774 0.504 0.664
512 DeLTa (GPT-4o) (Ours) 0.748 0.719 0.738 0.799 0.790 0.522 0.719 ($\uparrow$ 8.3%)
256 TabLLM (T0) 0.344 0.690 0.471 0.785 0.787 0.373 0.575
256 LIFT (GPT-3.5) 0.651 0.682 0.431 0.724 0.773 0.398 0.610
256 TP-BERTa (RoBERTa) 0.721 0.677 0.408 0.853 0.772 0.464 0.649
256 GTL (LLaMA2) - - - - - - -
256 LIFT-ICL (GPT-4o) - - - - - - -
256 P2T (GPT-4o) - - - - - - -
256 FeatLLM (GPT-4o) 0.729 0.658 0.459 0.740 0.772 0.473 0.639
256 DeLTa (GPT-4o) (Ours) 0.732 0.713 0.736 0.767 0.778 0.476 0.700 ($\uparrow$ 7.9%)
128 TabLLM (T0) 0.240 0.692 0.389 0.676 0.769 0.350 0.519
128 LIFT (GPT-3.5) 0.643 0.640 0.693 0.733 0.760 0.430 0.650
128 TP-BERTa (RoBERTa) 0.735 0.682 0.462 0.867 0.743 0.357 0.641
128 GTL (LLaMA2) - - - - - - -
128 LIFT-ICL (GPT-4o) - - - - - - -
128 P2T (GPT-4o) - - - - - - -
128 FeatLLM (GPT-4o) 0.721 0.648 0.568 0.698 0.768 0.474 0.646
128 DeLTa (GPT-4o) (Ours) 0.746 0.683 0.722 0.749 0.774 0.457 0.688 ($\uparrow$ 6.0%)
64 TabLLM (T0) 0.245 0.673 0.386 0.602 0.785 0.268 0.493
64 LIFT (GPT-3.5) 0.643 0.643 0.500 0.728 0.748 0.430 0.615
64 TP-BERTa (RoBERTa) 0.648 0.646 0.419 0.840 0.761 0.414 0.621
64 GTL (LLaMA2) 0.613 0.638 0.512 0.719 0.750 0.414 0.608
64 LIFT-ICL (GPT-4o) 0.348 0.538 0.326 0.469 0.613 0.251 0.424
64 P2T (GPT-4o) 0.368 0.542 0.334 0.569 0.634 0.243 0.448
64 FeatLLM (GPT-4o) 0.697 0.643 0.508 0.723 0.749 0.450 0.628
64 DeLTa (GPT-4o) (Ours) 0.732 0.663 0.615 0.733 0.758 0.457 0.660 ($\uparrow$ 5.0%)
16 TabLLM (T0) 0.240 0.665 0.365 0.489 0.501 0.215 0.413
16 LIFT (GPT-3.5) 0.670 0.608 0.408 0.583 0.705 0.363 0.556
16 TP-BERTa (RoBERTa) 0.659 0.609 0.355 0.625 0.734 0.394 0.563
16 GTL (LLaMA2) 0.604 0.622 0.324 0.628 0.746 0.387 0.552
16 LIFT-ICL (GPT-4o) 0.457 0.492 0.324 0.308 0.512 0.198 0.382
16 P2T (GPT-4o) 0.226 0.471 0.320 0.105 0.499 0.203 0.304
16 FeatLLM (GPT-4o) 0.588 0.626 0.450 0.629 0.728 0.399 0.570
16 DeLTa (GPT-4o) (Ours) 0.682 0.638 0.369 0.674 0.730 0.379 0.579 ($\uparrow$ 1.6%)
Table 11: Comparison with LLM-based methods on all regression datasets $\downarrow$. We also calculate the average relative improvement of DeLTa against the best baseline method ($\downarrow$ %). “# Num” indicates the number of training samples. This table serves as an extension of Table 9.
[…]
All TabLLM (T0) - - - - - - -
All LIFT (GPT-3.5) 0.379 0.983 0.629 0.902 0.518 0.328 0.623
All TP-BERTa (RoBERTa) 0.138 0.795 0.399 0.708 0.215 0.163 0.403
All GTL (LLaMA2) - - - - - - -
All LIFT-ICL (GPT-4o) - - - - - - -
All P2T (GPT-4o) - - - - - - -
All FeatLLM (GPT-4o) - - - - - - -
All DeLTa (GPT-4o) (Ours) 0.116 0.780 0.351 0.529 0.200 0.130 0.351 ($\downarrow$ 12.9%)
512 TabLLM (T0) - - - - - - -
512 LIFT (GPT-3.5) 0.432 1.167 0.752 1.385 0.589 0.368 0.782
512 TP-BERTa (RoBERTa) 0.186 0.837 0.528 0.760 0.383 0.280 0.496
512 GTL (LLaMA2) - - - - - - -
512 LIFT-ICL (GPT-4o) - - - - - - -
512 P2T (GPT-4o) - - - - - - -
512 FeatLLM (GPT-4o) - - - - - - -
512 DeLTa (GPT-4o) (Ours) 0.168 0.814 0.515 0.748 0.287 0.238 0.462 ($\downarrow$ 6.9%)
256 TabLLM (T0) - - - - - - -
256 LIFT (GPT-3.5) 0.324 1.034 0.683 1.294 0.578 0.379 0.715
256 TP-BERTa (RoBERTa) 0.359 0.863 0.557 0.798 0.545 0.368 0.582
256 GTL (LLaMA2) - - - - - - -
256 LIFT-ICL (GPT-4o) - - - - - - -
256 P2T (GPT-4o) - - - - - - -
256 FeatLLM (GPT-4o) - - - - - - -
256 DeLTa (GPT-4o) (Ours) 0.234 0.841 0.568 0.799 0.397 0.277 0.519 ($\downarrow$ 10.7%)
128 TabLLM (T0) - - - - - - -
128 LIFT (GPT-3.5) 0.290 1.070 0.710 1.300 0.660 0.430 0.743
128 TP-BERTa (RoBERTa) 0.431 0.903 0.597 0.880 0.654 0.390 0.642
128 GTL (LLaMA2) - - - - - - -
128 LIFT-ICL (GPT-4o) - - - - - - -
128 P2T (GPT-4o) - - - - - - -
128 FeatLLM (GPT-4o) - - - - - - -
128 DeLTa (GPT-4o) (Ours) 0.247 0.857 0.641 0.855 0.512 0.319 0.572 ($\downarrow$ 11.0%)
64 TabLLM (T0) - - - - - - -
64 LIFT (GPT-3.5) 0.610 1.050 0.810 1.320 0.700 0.470 0.827
64 TP-BERTa (RoBERTa) 0.645 0.928 0.704 0.920 0.717 0.445 0.727
64 GTL (LLaMA2) 0.663 1.190 0.834 1.189 0.724 0.483 0.847
64 LIFT-ICL (GPT-4o) 0.637 1.213 0.820 1.045 0.710 0.494 0.820
64 P2T (GPT-4o) 0.697 1.335 0.868 1.452 0.745 0.513 0.935
64 FeatLLM (GPT-4o) - - - - - - -
64 DeLTa (GPT-4o) (Ours) 0.382 0.888 0.701 0.919 0.634 0.383 0.651 ($\downarrow$ 10.4%)
16 TabLLM (T0) - - - - - - -
16 LIFT (GPT-3.5) 0.890 1.130 1.010 1.350 0.920 0.670 0.995
16 TP-BERTa (RoBERTa) 0.814 0.941 0.924 1.008 0.969 0.669 0.887
16 GTL (LLaMA2) 0.804 1.184 0.904 1.258 0.854 0.624 0.938
16 LIFT-ICL (GPT-4o) 0.815 1.256 1.074 1.172 0.977 0.855 1.025
16 P2T (GPT-4o) 0.795 1.390 0.961 1.438 0.844 0.824 1.042
16 FeatLLM (GPT-4o) - - - - - - -
16 DeLTa (GPT-4o) (Ours) 0.776 1.094 0.860 0.953 0.835 0.593 0.852 ($\downarrow$ 4.0%)
A.8 Full comparison results with conventional baseline methods
We also provide the full results of DeLTa against non LLM-based baseline methods in full-data regimes in Table 12 and Table 13.
Table 12: Comparison with non LLM-based methods on all classification datasets $\uparrow$. We also calculate the average relative improvement of DeLTa against the best baseline method ($\uparrow$ 3%).
[…]
KNN 0.7072 0.6429 0.7730 0.8903 0.8186 0.5168 0.7248
CART 0.7871 0.6829 0.7253 0.8900 0.8392 0.5642 0.7481
MLP 0.6882 0.7171 0.6974 0.9013 0.8523 0.7043 0.7601
RandomForest 0.8061 0.7371 0.7944 0.8915 0.8589 0.6588 0.7911
XGBoost 0.7605 0.7171 0.6102 0.9051 0.8709 0.7140 0.7630
Catboost 0.7757 0.7543 0.6036 0.9080 0.8726 0.7217 0.7727
ResNet 0.6464 0.7200 0.6332 0.9020 0.8512 0.7134 0.7444
FT-Transformer 0.7871 0.7229 0.7862 0.9030 0.8590 0.7272 0.7976
SAINT 0.7490 0.7486 0.7582 0.9051 0.8560 0.7050 0.7870
TabPFN (Version 2) 0.7985 0.7286 0.4507 0.9095 0.8611 0.7020 0.7417
ModernNCA 0.8099 0.7286 0.7368 0.9046 0.8702 0.7248 0.7958
DeLTa (Ours) 0.8289 0.7829 0.8355 0.9080 0.8677 0.7048 0.8213 ($\uparrow$ 3%)
Table 13: Comparison with non LLM-based methods on all regression datasets $\downarrow$. We also calculate the average relative improvement of DeLTa against the best baseline method ($\downarrow$ 2.4%).
[…]
KNN 0.2347 1.0975 0.5893 0.7970 0.4628 0.2135 0.5658
CART 0.2533 0.8196 0.7100 0.8433 0.6462 0.3504 0.6038
MLP 0.1318 0.8244 0.4628 0.5949 0.2072 0.3068 0.4213
RandomForest 0.1521 0.8093 0.5057 0.6692 0.3712 0.1626 0.4450
XGBoost 0.1285 0.7846 0.4037 0.6169 0.2291 0.1379 0.3835
Catboost 0.1210 0.7779 0.3723 0.5672 0.2024 0.1306 0.3619
ResNet 0.1362 0.8194 0.4434 0.5931 0.2083 0.2143 0.4025
FT-Transformer 0.1187 0.7822 0.3991 0.5851 0.2067 0.1680 0.3766
SAINT 0.1371 0.8276 0.4385 0.6007 0.2063 0.1396 0.3916
TabPFN (Version 2) 0.1375 0.7956 0.3440 0.5527 0.2001 0.1282 0.3597
ModernNCA 0.1242 0.7876 0.3656 0.6270 0.2027 0.1280 0.3725
DeLTa (Ours) 0.1156 0.7804 0.3507 0.5293 0.2000 0.1303 0.3511 ($\downarrow$ 2.4%)
A.9 Computational efficiency
Table 14: Runtime in seconds of DeLTa and other LLM-based baseline methods for the training and inference phase, conducted on Adult dataset. “# Num” indicates number of training samples. †These methods require querying LLMs for each test sample. * These methods do not require querying LLMs at inference phase.
# Num Stage TabLLM† LIFT† TP-BERTa† GTL† LIFT-ICL† P2T† FeatLLM* DeLTa* (Ours)
All Train 177371.37 153689.14 1319.35 N/A N/A N/A 1231.22 35.20
Inference 179.21 90149.48 1.58 N/A N/A N/A 0.14 0.09
64 Train 288.61 191.75 316.40 N/A N/A N/A 1047.55 23.09
Inference 170.42 89684.76 5.01 153501.02 176726.13 200258.24 0.14 0.08
A.10 Ablation results
Let $F(x)$ denote the well-trained ensemble of decision trees (Random Forest), and $\mathcal{R}$ the corresponding decision tree rule set. DeLTa first performs a rule refinement procedure via LLM to derive a new rule $r^{*}$, and subsequently applies an error correction mechanism to compute a sample-specific error correction vector based on $r^{*}$, which is then added to $F(x)$ to produce the final prediction. We further conduct ablation study to demonstrate the effectiveness of key components of DeLTa, as illustrated in Table 15 and Table 16, by comparing DeLTa against different groups of variants.
  • Variant A to B do not use the rule refinement process or the error correction process: (1) variant A, that is identical to the single decision tree (CART); (2) variant B, that is identical to the existing ensemble of decision trees $F(x)$ (Random Forest).
  • Variant C to D incorporate the rule refinement process and utilize the $r^{*}$ to predict labels, while removes the error correction process: (3) variant C, that treats the $r^{*}$ as a standalone decision tree by assigning the input $x$ to a specific leaf node; (4) variant D, that further appends $r^{*}$ to the existing ensemble of decision trees $F(x)$.
  • Variant E includes both the rule refinement process and the error correction process, and utilizes the $r^{*}$ to predict both labels and error correction vectors: (5) variant E, that uses $r^{*}$ to predict labels like variant D, and also uses $r^{*}$ to predict error correction vectors like DeLTa. The final output is the sum of the predicted label and the predicted error correction vector.
  • Variant F removes the rule refinement process, but remains the error correction process: (6) variant F, that replaces the LLM generated $r^{*}$ with $r\in\mathcal{R}$ derived from Random Forest to predict the error correction vector, and then adds it to the existing ensemble of decision trees $F(x)$ as the final prediction.
Table 15: Analysis on the effects of different components of DeLTa on classification tasks. “RF” corresponds to Random Forest.
[…]
A (CART) 0.787 0.683 0.725 0.890 0.839 0.564 0.748
B (RF) 0.806 0.737 0.794 0.892 0.859 0.659 0.791
C ✓(for label) 0.811 0.766 0.818 0.900 0.853 0.677 0.804
D ✓(for label) 0.812 0.771 0.819 0.899 0.856 0.671 0.805
E ✓(for residual and label) 0.816 0.779 0.829 0.895 0.858 0.679 0.809
F 0.810 0.726 0.778 0.892 0.842 0.613 0.777
DeLTa (Ours) ✓(for residual) 0.829 0.783 0.836 0.908 0.868 0.705 0.821
Table 16: Analysis on the effects of different components of DeLTa on regression tasks. “RF” corresponds to Random Forest.
[…]
A (CART) 0.253 0.820 0.710 0.843 0.646 0.350 0.604
B (RF) 0.152 0.809 0.506 0.669 0.371 0.163 0.445
C ✓(for label) 0.136 0.808 0.402 0.607 0.272 0.139 0.394
D ✓(for label) 0.144 0.800 0.387 0.583 0.258 0.138 0.385
E ✓(for residual and label) 0.125 0.799 0.379 0.552 0.251 0.142 0.375
F 0.218 0.819 0.544 0.811 0.619 0.324 0.556
DeLTa (Ours) ✓(for residual) 0.116 0.780 0.351 0.529 0.200 0.130 0.351
In the Table 4 in the original paper, DeLTa w/o “RR” corresponds to the variant F, and DeLTa w/o “EC” corresponds to the variant C. Here, “RR” denotes the rule refinement process, and “EC” denotes the error correction process. The results show that the decision tree rule refinement via LLM is important. Directly using the refined rule $r^{*}$ to predict labels could also enhance the performance (variant C and D). And the error correction strategy could further enhance tabular prediction performance (variant E and DeLTa). However, without $r^{*}$, the effect of error correction is limited (variant F). Therefore, the removal of any of the components degrades the performance of DeLTa.
A.11 Hyper parameter sensitivity analysis
We incorporate the sensitivity analysis for the number of LLM querying times for DeLTa in Fig. 7. It is reasonable for us to set the querying times to 10 by default.
Figure 7: The effect of the number of LLM queries on performance averaged over all datasets.
A.12 Additional analysis
Fig. 8 shows the label predictions of ours: $F(x)+\Delta_{x}$ and ours w/o $\Delta_{x}$: $F(x)$. We can observe that two categories of samples are overlapped in some regions. $F(x)$ struggles to distinguish them, but DeLTa could correct the prediction errors of $F(x)$ to enhance the prediction for such complex data patterns.
(a) Ground Truth
(b) Predicted by RandomForest
(c) Predicted by ours
Figure 8: T-SNE visualization of label prediction of DeLTa w/ and w/o error correction vector $\Delta_{x}$ (i.e., Random Forest) on BA dataset.
We compute the intra-node sample distance averaged over all leaf nodes partitioned by refined rule $r^{*}$ and random forest rules, respectively. The results in Table 17 show that LLM could generate a better rule, where samples grouped within the same leaf node exhibit greater statistical similarity.
Table 17: Average intra-node distance comparison.
BL Credit Car Bank AD JA Average
RandomForest Rule 2.8520 6.0938 2.6609 5.0186 4.7781 9.9177 5.2202
Refined Rule 1.3551 5.0503 2.5059 4.7078 4.2950 9.4011 4.5525 ($\downarrow$ 12.79%)
We also provide an example of the original decision tree rule set derived from Random Forest in Fig. 9, along with the corresponding refined rule generated by the LLM in Fig. 10. This example is taken from the low-data regime with 16 training samples on the Credit dataset.
Figure 9: An example of three tree rules obtained by the Random Forest.
Figure 10: An example of LLM’s response for rule refinement.
How an LLM refines decision tree rules.
(i) The intuition. Traditional methods like CART use information gain to make a series of locally optimal decisions, allowing a data-driven algorithm to find the best feature splits. This can lead to rules that are brittle or overfit to the specific data sample used for that tree. DeLTa, instead, utilizes the LLM to analyze diverse rules from a Random Forest and performs a logic-based reasoning task: (a) It recognizes patterns across multiple “expert” rules to identify features and decision paths that are consistently important. (b) It abstracts the core logic splits from individual trees. (c) It synthesizes a new, globally coherent rule that represents a consensus of the most robust patterns.
(ii) A concrete example of a better rule. As shown in Fig. 9 and Fig. 10, we provide a set of decision tree rules from Random Forest and refined rule by the LLM respectively. This is from the Credit dataset, where each sample (person) is classified as good or bad credit risks according to the set of features. To provide deeper insight into how the refined rule improves upon the original ones, we also provide the average SHAP values [lundberg2017unified] of features under the Random Forest and our method respectively in Table 18. SHAP assigns each feature an importance value (higher is more important) for a particular prediction, thus it serves an unified framework for interpreting predictions.
The results show that, after LLM refinement, the SHAP importance of features such as “credit_history” (feature_16) and “employment” (feature_10) are increased, which aligns with the financial expertise. Let us take a closer look into Fig. 9 and Fig. 10, for “credit_history” (feature_16), this feature appears near the top of the second tree, indicating its potential significance. However, only 1 out of 3 decision trees in the original Random Forest rule set explicitly splits on this feature, which may dilute its influence in the ensemble prediction. For “employment” (feature_10), although it appears in 2 out of 3 rules, its splits occur late in the tree, leading to a low overall contribution in prediction. In contrast, the LLM-refined rule elevates the prominence of these features.
(iii) Why do traditional techniques not capture such a rule? More broadly, a single decision tree with information gain has the risk of overfitting. Although Random Forest ensembles the outputs from all tree rules, the inherent relationships and interactions among these rules are ignored. As the number of trees increases, analyzing these independent rules is gradually becoming more and more difficult, let alone utilizing these rules to partition feature space. To this end, we propose to leverage LLMs to analyze and summarize these rules into a refined rule due to the powerful logical reasoning ability of LLM. This enables the generation of improved partitioning strategies over the feature space, thereby promoting statistical coherence among samples assigned to the same leaf node. To verify it, we compute the intra-node sample distance over all leaf nodes partitioned by LLM refined rule $r^{*}$, and observe that the distance of $r^{*}$ is lower than original rule $r\in\mathcal{R}$ from Random Forest, as illustrated in Fig. 2 and Table 17.
Table 18: SHAP values of different features.
Original property_magnitude other_payment_plans checking_status purpose housing foreign_worker job credit_history savings_status personal_status
0.1487 0.0471 0.0410 0.0355 0.0326 0.0251 0.0213 0.0189 0.0105 0.0070
Ours property_magnitude credit_history checking_status purpose employment other_payment_plans foreign_worker housing savings_status job
0.1055 0.0436 0.0342 0.0174 0.0155 0.0144 0.0132 0.0106 0.0096 0.0095
Assessing the inherent quality of LLM-rewritten rules. We assess this in a principled way from three perspectives: (i) SHAP value analysis, (ii) intra-node distance comparison and (iii) the transparency of the LLM’s process, where the LLM provides a human-readable rationale for its actions, which allows us to inspect and understand its reasoning process.
Comparison on additional datasets. We have include additional datasets from the WhyTrees [grinsztajn2022tree] benchmark. Specifically, we selected a subset of representative datasets that span different scales and feature characteristics, including the large-scale, high-dimensional “Year” dataset, to evaluate DeLTa’s scalability and generalizability. The updated results, summarized in Table 19 and Table 20, further demonstrate DeLTa’s strong and consistent performance across a more diverse and standardized set of tabular tasks.
Table 19: Additional tabular dataset properties. “#objects” indicates the number of samples in the dataset. “#num. features” indicates the number of numerical features, and “#cat. features” indicates the number of categorical features.
Dataset #objects #num. features #cat. features metric #classes
rl 4970 5 7 Acc. 2
GiveMeSomeCredit 16714 10 0 Acc. 2
Year 515345 90 0 NRMSE
Table 20: Comparison with baseline methods on additional datasets. “(L)” and “(T)” indicates correspond to LLM-based and traditional methods, respectively.
Dataset TabLLM (L) FeatLLM (L) LIFT (L) TP-BERTa (L) KNN (T) CART (T) MLP (T) RandomForest (T) XGBoost (T) CatBoost (T) FT-Transformer (T) DeLTa (L)
rl $\uparrow$ 0.7716 0.7243 0.6349 0.7394 0.6258 0.6308 0.6630 0.7022 0.7726 0.7746 0.7072 0.7897
GiveMeSomeCredit $\uparrow$ 0.7765 0.7137 0.6534 0.7810 0.6063 0.7514 0.7469 0.7768 0.7789 0.7792 0.7610 0.7816
Year $\downarrow$ - - 1.0540 0.8258 1.0459 0.9374 0.8134 0.8547 0.8376 0.8157 0.8090 0.8015
A.13 Discussion
The existing LLM-based methods for tabular prediction suffer from two key inherent issues: (i) data perspective: existing data serialization methods lack universal applicability for structured tabular data, and may pose privacy risks through direct textual exposure, and (ii) model perspective: LLM fine-tuning methods struggle with tabular data, and in-context learning scalability is bottle-necked by input length constraints (suitable for few-shot learning). The proposed DeLTa could well solve the aforementioned challenges: Solving the data issue: Unlike serialization methods that convert each sample into unnatural text formats, decision tree rules are composed of simple comparisons between feature values and thresholds, forming logical, interpretable structures that can be naturally expressed in text without relying on semantic columns names. In addition, decision tree rules represent global feature space partitioning rule rather than individual samples, which helps mitigate privacy concerns by avoiding exposure of sample-level information. Solving the model issue: Moreover, the powerful reasoning ability of LLMs can be leveraged to redesign decision tree rules and help trees with aggregating their decisions, rather than directly using LLMs to generate label predictions. Notably, DeLTa avoids serializing tabular data into natural language format, and does not require additional domain-specific expertise or semantic information, such as explicit feature names and detailed task background knowledge. Furthermore, DeLTa can be applied in full data learning setting without LLM fine-tuning.
Overall, DeLTa achieves the highest average performance in all settings, including classification and regression, in both full and low-data regimes. In addition, DeLTa is computationally efficient in both training and inference stages. This efficiency stems from key properties: (i) DeLTa utilizes the reason ability of LLM without fine-tuning LLM, (ii) DeLTa avoids querying LLMs to generate predictions for individual samples. Instead, it only needs to query LLM via API to generate one refined decision tree rule for one dataset, enabling significantly more efficient use of LLM resources.
A.14 Limitation
Although the primary focus of this work is tabular data, the flexibility of DeLTa opens up exciting possibilities for extensions. For example, DeLTa can be extended to other domains such as time series, graph and so on. We leave the extensions of DeLTa as a future work. While DeLTa avoids tabular data serialization, it requires training decision trees and extracting corresponding decision rules. However, single decision tree may be prone to overfitting, potentially resulting in suboptimal rules and limited performance improvements in our framework. To mitigate this issue, we adopt Random Forest to generate a diverse set of decision tree rules. This strategy is not only effective but also easy to implement using standard libraries such as scikit-learn, which conveniently supports both model training and rule extraction.
Another limitation of DeLTa is that its performance advantage over non-LLM baselines narrows with increasing dataset sizes. Upon deeper analysis, we have found that DeLTa’s performance advantage is more pronounced when the dataset size is relatively small, compared to non-LLM baselines. Conversely, the performance advantage of data-intensive baseline methods, such as neural networks, tends to increase with larger dataset sizes. This pattern may be attributed to two main factors: (i) Nature of Base Rules: Our method refines the base rules extracted from Random Forest (RF). While RF is effective due to its ensemble nature, its trees are independently trained without joint optimization. As the data volume increases, this independent training may limit RF’s ability to fully leverage large-scale data. Consequently, the quality of the base rules may plateau with increasing data, which also affects the upper bound of the refined rules. (ii) Comparison to Neural Networks (NNs): In contrast, neural networks are typically data-intensive models that require large amounts of training data to perform well and generalize effectively. Their performance tends to scale better with larger datasets, as they can learn complex feature interactions.
A.15 Future direction
A key direction involves developing more scalable base rule extraction mechanisms. Exploring ensembles of jointly optimized trees or hybrid models that better leverage large-scale data could raise the performance ceiling for refinement. Furthermore, integrating the rule-based paradigm with data-intensive architectures—for instance, by using neural networks to generate or pre-select candidate rules, or by designing LLM-driven fine-tuning that adapts to dataset scale—presents a promising avenue to combine the interpretability of rules with the scalability of neural models.
A.16 Broader impacts
DeLTa advances the integration of LLMs with structured tabular data by introducing a novel way that leverages logical decision tree rules as intermediaries. By representing the entire dataset through decision tree rules, DeLTa avoids the need for sample-level serialization. This makes it particularly well-suited for structured tabular data with heterogeneous features, in contrast to the unstructured nature of textual data. DeLTa also offers practical benefits in terms of data privacy. It could help mitigate privacy concerns by avoiding exposure of sample-level information, which is useful in privacy-critical domains such as healthcare and finance. Furthermore, DeLTa can be applied in full data learning setting without LLM fine-tuning, thereby offering a cost-efficient and scalable solution for integrating LLMs into tabular learning tasks.